### v4

# Differential Confounding Privacy and Inverse Composition

Abstract
Differential privacy (DP) has become the gold standard for privacy-preserving data analysis, but its applicability can be limited in scenarios involving complex dependencies between sensitive information and datasets. To address this, we introduce differential confounding privacy (DCP), a specialized form of the Pufferfish privacy (PP) framework that generalizes DP by accounting for broader relationships between sensitive information and datasets. DCP adopts the $(\epsilon,\delta)$-indistinguishability framework to quantify privacy loss. We show that while DCP mechanisms retain privacy guarantees under composition, they lack the graceful compositional properties of DP. To overcome this, we propose an Inverse Composition (IC) framework, where a leader-follower model optimally designs a privacy strategy to achieve target guarantees without relying on worst-case privacy proofs, such as sensitivity calculation. Experimental results validate IC’s effectiveness in managing privacy budgets and ensuring rigorous privacy guarantees under composition.
I Introduction
Information privacy is a critical challenge in the digital age, driven by the widespread use of data-driven technologies and personal data. Addressing this requires robust, interpretable, and quantifiable privacy protection techniques. The problem can be framed as follows: Let $X$ be a dataset processed by a (possibly randomized) mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$, producing an output $Y=\mathcal{M}(X)$ that is publicly observed, where $\mathcal{X}$ and $\mathcal{Y}$ are the input and output spaces, respectively. The dataset $X$ encodes information about sensitive information (or the secret) $S$. The goal of privacy protection is to quantify how much an attacker can infer about the secret $S$ from the output $Y$.
Differential privacy (DP) [1] has become the gold standard for privacy-preserving data analysis. In DP, the secret $S$ is an integral part of a dataset, encompassing attributes, entries, values, or membership. DP ensures that any two adjacent datasets $x,x^{\prime}\in\mathcal{X}$, differing by a single entry ($x\simeq x^{\prime}$), produce nearly indistinguishable outputs, limiting the information inferable about the secret. A randomized mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-differentially private ($(\epsilon,\delta)$-DP) with $\epsilon\geq 0$ and $\delta\in[0,1]$ if it is $(\epsilon,\delta)$-indistinguishable ($(\epsilon,\delta)\textup{-}\mathtt{ind}$) for all $x\simeq x^{\prime}$:
$\sup_{\mathcal{W}\subset\mathcal{Y}}\left(\textbf{Pr}\left[\mathcal{M}(x)\in\mathcal{W}\right]-e^{\epsilon}\textbf{Pr}\left[\mathcal{M}(x^{\prime})\in\mathcal{W}\right]\right)\leq\delta.$ (1)
Pufferfish privacy (PP) [2] generalizes differential privacy (DP) by extending the adjacency concept from datasets to distributions over datasets, conditioned on the adversary’s prior knowledge [3]. This allows PP to incorporate domain-specific assumptions and protect customizable sensitive properties of the data—extending beyond individual data points or records. Appendix -B shows the formal definition of PP.
In DP, the secret and the dataset enter the $(\epsilon,\delta)$-indistinguishability inequality through a deterministic one-to-one relationship: once the dataset is fixed, the secret (i.e., data point) is uniquely determined, and adjacent datasets correspond to distinct secret values. Hence, the randomness in the relationship between the secret and the mechanism output arises solely from the randomness of the mechanism itself. However, in real-world systems such as multi-agent networks, machine learning pipelines, or distributed systems, sensitive information often arises from complex interactions and dependencies that go beyond individual records, data attributes, or features that are observable or directly inferable from the dataset (e.g., an individual’s activity at a specific time). These systems involve derived knowledge, intermediate information, or correlations generated within internal mechanisms. Here are a few illustrative examples (see Appendix -D for more detailed examples).
Example: Online Learning and Personalized Recommendation Systems
Online platforms use user interactions $X$ (e.g., clicks, watch history) to infer secrets $S$, such as latent preferences or user profiles. The platform updates $S$ based on $X$, and $S$ shapes future recommendations, influencing user behavior. Over time, $S$ and $X$ evolve together in a feedback loop ($S\leftrightarrow X$). Unlike static datasets, $S$ is dynamically updated and not directly present in $X$.
Example: Medical Diagnosis and Screening Pipelines
In healthcare, tests and screenings generate datasets $X$ (e.g., lab results, imaging data). The secret $S$ is a diagnosis or health label inferred from $X$, such as “high risk of condition $S$.” The causal flow starts with raw data $X$, which is analyzed to derive $S$. Feedback loops occur when a tentative diagnosis triggers further tests, creating new data that updates $X$. Since $S$ depends on multiple variables in $X$ and evolves dynamically, $S\not\subseteq X$.
To capture such general relationships between $S$ and $X$, we introduce differential confounding privacy (DCP), a specialized form of the PP framework, to quantify the privacy leakage of the secret $S$ from the output $Y$. DCP adopts the $(\epsilon,\delta)$ scheme of DP and PP to parameterize probabilistic (near) indistinguishability.
Composition, the ability to run multiple private algorithms on the same dataset while preserving privacy, is a fundamental property of effective privacy frameworks. Graceful composition properties are key to the success of DP [4] and are essential for the practicality and scalability of any useful privacy framework. We show that the composition of DCP mechanisms remains DCP but lacks the graceful composition properties of DP even for independent mechanisms. Specifically, the optimal composition bound for DP [5, 6] underestimates the cumulative privacy loss in DCP, and the basic composition property [7, 5] is generally not satisfied. This presents challenges in determining optimal or conservative privacy bounds for DCP under composition.
To tackle these challenges, we introduce the inverse composition (IC) framework, which reimagines the approach to privacy composition. In contrast to traditional methods that depend on worst-case guarantees, such as sensitivity calculations (generally NP-hard [8]) and combinatorial composition analysis (generally #P-hard [6]), IC adopts a leader-follower model to strategically design privacy strategies. The leader’s privacy strategy elicits an optimal response from the follower, ensuring that privacy guarantees are met without explicitly constraining the privacy strategy or the mechanism. IC provides a systematic method to achieve $(\epsilon,\delta)$-DCP guarantees under composition, offering a flexible framework for managing privacy budgets and addressing the complexities of DCP.
Organization: Section II defines the DCP framework. Section III analyzes the composition of DCP mechanisms and their limitations compared to DP. Section IV presents the inverse composition framework, and Section V provides numerical validation. Section VI concludes the paper. Technical details and proofs are in the appendices of the full version [9].
I-A Related Work
Differential confounding privacy (DCP) builds on the general Pufferfish privacy (PP) framework [2], which extends the deterministic, one-to-one relationship between secrets and datasets assumed in Differential privacy (DP) [1] to support more general and expressive privacy specifications. Despite its expressive power, a key challenge in the PP framework is the scarcity of practical and broadly applicable mechanisms. Notable instantiations include Blowfish privacy [10], Wasserstein mechanisms for correlated data [11], attribute privacy [12], and mutual-information-based PP [13].
Composition of PP PP does not always compose [2, 11]. [2] introduce a sufficient condition, universally composable evolution scenarios (UCES), for linear self-composition, analogous to DP’s basic composition property [1]. UCES assumes a deterministic one-to-one mapping between a secret and a dataset, leveraging DP’s composition benefits by eliminating randomness in this relationship, which corresponds to an extremely confident attacker. However, in CP, where causation may flow from datasets to secrets, UCES is often infeasible. This complexity arises when another internal mechanism controls the state-dataset relationship, preventing the elimination of randomness.
Classic Composition of DP DP’s composition property underpins its success. The advanced composition of $(\epsilon,\delta)$-DP [14] tightens cumulative privacy bounds by considering probabilistic behavior. [5] derive optimal composition for homogeneous $(\epsilon,\delta)$-DP mechanisms, while [6] show that finding tight bounds for heterogeneous mechanisms is #$\#P$-complete. Recent works on interactive mechanisms confirm that they compose comparably to standard DP mechanisms [15, 16, 17].
Fine-Grained Composition of DP Functional approaches like Rényi DP (RDP) [18], privacy profiles [19], $f$-DP [20], and privacy loss distribution (PLD) [21, 22] provide fine-grained characterizations of privacy guarantees. These methods compose by aggregating functional terms, enabling mechanism-specific analyses, tighter privacy bounds, better privacy-utility trade-offs, and efficient privacy accounting [23, 24, 19, 25, 22, 26, 27, 28]. Conversions to $(\epsilon,\delta)$-DP are achieved through mechanism standardization, but lossless conversions avoiding #$\#P$-completeness under composition remain an open problem.
II Differential Confounding Privacy
We formally define differential confounding privacy (DCP) as follows. Let $\{\theta,\mathcal{G}\}$ describe the relationship between the secret $S$ and the dataset $X$, where:
  • $\theta\in\Theta\subseteq\Delta(\mathcal{S}\times\mathcal{X})$ determines a joint density $P_{\theta}(s,x)$, where $\Theta$ is a class of joint distributions over $S$ and $X$.
  • $\mathcal{G}:\mathcal{X}\to\mathcal{S}$ is a randomized mapping such that $P_{\theta}(s,x)>0$ if and only if $\textbf{Pr}[\mathcal{G}(x)=s]>0$.
In addition, let $P_{\theta}(s)$ and $P_{\theta}(x|s)$ be the marginal and conditional densities derived from $P_{\theta}(s,x)$. In addition, let $\mathtt{D}(\cdot|\theta):\mathcal{S}\times\mathcal{S}\to\mathbb{R}$ be a metric satisfying $\mathtt{D}(s,s^{\prime}|\theta)\leq\mathtt{d}$ for a given $\mathtt{d}\geq 0$ for any $s,s^{\prime}\in\mathcal{S}$ with $P_{\theta}(s)>0$ and $P_{\theta}(s^{\prime})>0$, which defines when two secrets $s$ and $s^{\prime}$ are considered adjacent. Let $\mathcal{Q}$ be the set of such adjacent secrets.
DCP aims to protect the privacy of the secret $S$ by ensuring that any adjacent secret pair from the output of a randomized mechanism $\mathcal{M}(\gamma)$, where $\gamma:\mathcal{X}\to\Delta(\mathcal{Y})$ denotes the probability density function of the underlying distribution of the mechanism. The corresponding mapping $\mathcal{M}(\cdot;\gamma):\mathcal{X}\to\mathcal{Y}$ takes an input dataset $x\in\mathcal{X}$ (but not the secret $S$) and produces an output $y=\mathcal{M}(x;\gamma)\in\mathcal{Y}$. For any measurable set $\mathcal{W}\subseteq\mathcal{Y}$, the probability is defined as:
$\textbf{Pr}^{\theta,s}_{\gamma}\left[Y\in\mathcal{W}\right]\equiv\int_{x\in\mathcal{X}}\int_{y\in\mathcal{W}}\gamma\left(y\middle|x\right)P_{\theta}(x|s)\,dy\,dx.$ (2)
For discrete spaces, the density functions are replaced by probability mass functions, and integrals are replaced by summations. This work focuses on the continuous case. For simplicity, we may omit $\theta$ in the notation $\textbf{Pr}^{\theta,s}_{\gamma}$ when $\theta$ is fixed and does not vary in the discussion.
Definition 1 ($(\epsilon,\delta,\theta,\mathcal{G})$-Differential Confounding Privacy)
A randomized mechanism $\mathcal{M}(\gamma):\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta,\theta,\mathcal{G})$-differentially confounding private ($(\epsilon,\delta,\theta,\mathcal{G})$-DCP) with $\epsilon\geq 0$ and $\delta\in[0,1]$ if $\mathcal{M}$ satisfies $(\epsilon,\delta,\theta,\mathcal{G})$-indistinguishability ($(\epsilon,\delta,\theta,\mathcal{G})\textup{-}\mathtt{ind}$):
$\sup_{(s,s^{\prime})\in\mathcal{Q}}\sup_{\mathcal{W}\subset\mathcal{Y}}\left(\textbf{Pr}^{s}_{\gamma}\left[Y\in\mathcal{W}\right]-e^{\epsilon}\textbf{Pr}^{s^{\prime}}_{\gamma}\left[Y\in\mathcal{W}\right]\right)\leq\delta.$ (3)
DCP ensures that any two adjacent secrets $(s,s^{\prime})\in\mathcal{Q}$ link to nearly indistinguishable outputs, thereby limiting the information that can be inferred about the secret. Specifically, the $(\epsilon,\delta,\theta,\mathcal{G})\textup{-}\mathtt{ind}$ condition bounds the difference in probabilities of $\mathcal{M}$ producing an output in any measurable set $\mathcal{W}\subset\mathcal{Y}$ for $s$ and $s^{\prime}$, up to an additive factor $\delta$ and multiplicative factor $e^{\epsilon}$. Smaller $\epsilon$ and $\delta$ imply stronger privacy guarantees. For simplicity, we omit $\mathtt{D}$ and $\mathtt{d}$ in the notation and refer to the property as $(\epsilon,\delta)$-DCP and $(\epsilon,\delta)\textup{-}\mathtt{ind}$.
Example: Privacy Risks in Navigation Analytics. Let $\mathbf{x}$ be the data from a connected vehicle, including GPS logs, driving speed, acceleration patterns, daily mileage, time-of-day usage, and onboard sensor readouts. Let $\mathbf{s}$ be the driver’s hidden risk profile (e.g., “high likelihood of accident or reckless driving”), inferred by an insurer or telematics service using a machine-learning model. A mechanism $\mathcal{M}(\gamma)$ performs a navigation analytics service that uses the raw data $\mathbf{x}$ to identify traffic congestion hotspots and suggest better route planning for municipal authorities or map providers. The output $\mathbf{y}=\mathcal{M}(\mathbf{x};\gamma)$ might include publicly accessible heatmaps of traffic speeds, congestion levels, and dangerous intersections across the city. Even without direct access to $\mathbf{x}$, an insurance company could leverage $\mathbf{y}$ to infer details about individual drivers. For instance, granular heatmaps could reveal where and when specific vehicles frequently drive or expose unique behaviors like repeated harsh braking on certain routes. By cross-referencing this information with partial records (e.g., billing data) and applying advanced machine-learning models, the insurer could refine risk scores or reconstruct the driver’s hidden risk profile $\mathbf{s}$.
In this example, the secret $s$ is inferred information derived from $x$. Consequently, DP becomes inapplicable, as it is designed to protect individual input data records rather than properties inferred from them.
III Characterizing the Composition
In this section, we characterized the composition properties of DCP mechanisms using copula and privacy loss random variables. We demonstrate that, unlike DP, DCP does not always exhibit composition as gracefully as DP when the same dataset (or overlapping datasets [14]) linked to the secret are used across multiple private data processing mechanisms. The detailed version of this section is provided in Appendix -C.
Composition In the DCP framework, as more computations are performed on the dataset $x$, it is important to understand how the privacy of the secret $s=\mathcal{G}(x)$ degrades under composition. Consider $k$ independent mechanisms $\mathcal{M}_{1}(\gamma_{1}),\mathcal{M}_{2}(\gamma_{2}),\dots,\mathcal{M}_{k}(\gamma_{k})$, where each $\mathcal{M}_{i}(\cdot;\gamma_{i}):\mathcal{X}\to\mathcal{Y}_{i}$ operates with a density function $\gamma_{i}$ and output space $\mathcal{Y}_{i}$. Let $\vec{\gamma}\equiv(\gamma_{i})_{i=1}^{k}$ denote the collection of density functions, $\vec{Y}\equiv(Y_{i})_{i=1}^{k}$ the random variables corresponding to the outputs of the mechanisms, and $\vec{y}\equiv(y_{i})_{i=1}^{k}\in\vec{\mathcal{Y}}\equiv\prod_{i=1}^{k}\mathcal{Y}_{i}$. The composition $\mathcal{M}(\cdot;\vec{\gamma}):\mathcal{X}\to\vec{\mathcal{Y}}$ is defined as:
$\mathcal{M}(x;\vec{\gamma})=\left(\mathcal{M}(x;\gamma_{1}),\dots,\mathcal{M}(x;\gamma_{k})\right).$
For any $\vec{\mathcal{W}}\subseteq\vec{\mathcal{Y}}$, we define:
$\textbf{Pr}^{s}_{\vec{\gamma}}\left[\vec{Y}\in\vec{\mathcal{W}}\right]\equiv\int\limits_{x\in\mathcal{X}}\int\limits_{\vec{y}\in\vec{\mathcal{W}}}\prod_{i=1}^{k}\gamma_{i}(y_{i}\mid x)P_{\theta}(x|s)\,d\vec{y}\,dx.$ (4)
The ($k$-fold) composition $\mathcal{M}(\vec{\gamma})$ satisfies $(\epsilon_{g},\delta_{g})$-DCP for some $\epsilon_{g}\geq 0$ and $\delta_{g}\in[0,1]$ if,
$\sup_{(s,s^{\prime})\in\mathcal{Q}}\sup_{\vec{\mathcal{W}}\subset\vec{\mathcal{Y}}}\left(\textbf{Pr}^{s}_{\vec{\gamma}}\left[\vec{Y}\in\vec{\mathcal{W}}\right]-e^{\epsilon_{g}}\textbf{Pr}^{s^{\prime}}_{\vec{\gamma}}\left[\vec{Y}\in\vec{\mathcal{W}}\right]\right)\leq\delta_{g}.$
Given a $\delta_{g}$, the tightest or optimal value of $\epsilon_{g}$ is defined as
$\textup{OPT}\left(\mathcal{M}(\vec{\gamma}),\delta_{g}\right)\equiv\inf\left\{\epsilon^{\prime}\geq 0\mid\mathcal{M}(\vec{\gamma})\textup{ is $(\epsilon^{\prime},\delta_{g})\textup{-}\mathtt{ind}$}\right\}.$ (5)
Effective Mechanism For each $\gamma_{i}:\mathcal{X}\to\Delta(\mathcal{Y}_{i})$, let $\psi_{i}:\mathcal{S}\to\Delta(\mathcal{Y}_{i})$ denote the underlying density function of $\textbf{Pr}^{s}_{\gamma}[\cdot]$ given by Equation (2). Specifically,
[…]
Additionally, let $\mathcal{N}_{i}(\cdot;\psi_{i}):\mathcal{S}\to\mathcal{Y}_{i}$ denote the corresponding effective mechanism for all $i\in[k]$. Since $\textbf{Pr}^{s}_{\psi_{i}}[\cdot]=\textbf{Pr}^{s}_{\gamma_{i}}[\cdot]$ for all $i$, the mechanism $\mathcal{N}_{i}(\psi_{i})$ is $(\epsilon_{i},\delta_{i})\textup{-}\mathtt{ind}$ if and only if $\mathcal{M}_{i}(s;\gamma_{i})$ is $(\epsilon_{i},\delta_{i})$-DCP. Let $\Psi_{i}$ denote the cumulative distribution function (CDF) associated with the density $\psi_{i}$ of each effective mechanism $\mathcal{N}_{i}$. The composition of the effective mechanisms $\mathcal{N}(\cdot;\vec{\psi}):\mathcal{S}\to\vec{\mathcal{Y}}$, defined as $\mathcal{N}(s;\vec{\psi})=\left(\mathcal{N}_{1}(s;\psi_{1}),\dots,\mathcal{N}_{k}(s;\psi_{k})\right)$.
If we consider the secret as a single-point dataset (see Appendix -E for discussion of signle-point DP), $\textup{OPT}\left(\mathcal{N}(s;\vec{\psi}),\delta_{g}\right)$ can be calculated according to the optimal composition theory of DP [6]. However, it is generally a #P-complete problem [6].
The concept of effective mechanisms is introduced to illustrate the connection and the difference between the composition properties of DP and those of DCP, in terms of privacy loss random variable.
Privacy Loss Random Variable Privacy loss under the $(\epsilon,\delta)$ framework can be quantified through the privacy loss random variable (PLRV). Let $\mathtt{p}$ and $\mathtt{q}$ be two probability density functions on the output space $\mathcal{Y}$. Define the PLRV as
$\textbf{L}_{\mathtt{p},\mathtt{q}}(y)=\log\frac{\mathtt{p}(y)}{\mathtt{q}(y)}.$
The PLRV, $\textbf{L}_{\mathtt{p},\mathtt{q}}(Y)$ for $Y\sim\mathtt{p}$, captures the distinguishability between outputs under $\mathtt{p}$ and $\mathtt{q}$. Let $\psi^{s}_{i}(\cdot)=\psi_{i}(\cdot|s)\in\mathcal{Y}_{i}$. Then, for any $(s_{0},s_{1})\in\mathcal{Q}$, the individual PLRV of each mechanism $\mathcal{M}_{i}(\gamma_{i})$ is given by $\textbf{L}_{\psi^{s_{0}}_{i},\psi^{s_{1}}_{i}}(Y_{i})$ with $Y_{i}\sim\psi^{s_{0}}_{i}$.
If we consider a deterministic, one-to-one relationship between $S$ and $X$, DCP becomes DP. The PLRV of the composition $\mathcal{M}(\vec{\gamma})$ coincides with the PLRV of the composition of efficient mechanisms $\mathcal{N}(\vec{\psi})$. That is,
$\textbf{L}^{\textup{id}}_{\vec{\psi}_{s_{0}},\vec{\psi}_{s_{1}}}(\vec{Y})=\sum\nolimits_{i=1}^{k}\textbf{L}_{\psi_{i}^{s_{0}},\psi_{i}^{s_{1}}}(Y_{i}),$ (6)
where $\vec{\psi}_{s}(\cdot)=\vec{\psi}(\cdot|s)=\prod^{k}_{i=1}\psi_{i}(\cdot|s)$. The fact that privacy loss can be either positive or negative introduces cancellations under composition, which are generally challenging to track.
As we will show shortly using copulas, under a general $(\theta,\mathcal{G})$, the PLRV of $\mathcal{M}(\vec{\gamma})$ differs from $\textbf{L}^{\textup{id}}{\vec{\psi}{s_{0}},\vec{\psi}{s{1}}}(\vec{Y})$, even when the mechanisms $\{\mathcal{M}_{i}(\gamma_{i})\}^{k}_{i=1}$ are independent.
Copula For any $s\in\mathcal{S}$, let $B(\cdot|s)$ denote the cumulative distribution function (CDF) and $b(\cdot|s)$ the density function of $\textbf{Pr}^{s}_{\vec{\gamma}}$ as defined in Equation (10) of Appendix -C. To simplify, we assume the outputs $y_{i}\in\mathcal{Y}_{i}$ are univariate for all $i\in[k]$. By Sklar’s theorem [29], there exists a copula $C$ for every $s\in\mathcal{S}$ such that:
$B\left(Y_{1},\dots,Y_{k}\mid s\right)=C\left(\Psi_{1}\left(Y_{1}\mid s\right),\dots,\Psi_{k}\left(Y_{k}\mid s\right)\mid s\right),$
where $C:[0,1]^{k}\to[0,1]$ is a multivariate CDF with uniform univariate marginals:
$C(u_{1},\dots,u_{k})=\textbf{Pr}\left[U_{1}\leq u_{1},\dots,U_{k}\leq u_{k}\right],$
and $U_{i}\in[0,1]$ is a standard uniform random variable $\forall i\in[k]$.
Let $c$ denote the copula density function of $C$. Then, the density function $b$ of $B$ can be expressed as:
$b_{s}\left(\vec{Y}\right)=c\left(\Psi_{1}\left(Y_{1}\mid s\right),\dots,\Psi_{k}\left(Y_{k}\mid s\right)\mid s\right)\prod_{i=1}^{k}\psi_{i}\left(Y_{i}\mid s\right).$
Given $\{\theta,\mathcal{G}\}$, the PLRV of the composition $\mathcal{M}(\vec{\gamma})$ is expressed as:
$\textbf{L}_{b_{s_{0}},b_{s_{1}}}(\vec{Y})=\textbf{L}^{\mathcal{G}}_{c_{s_{0}},c_{s_{1}}}(\vec{Y})+\textbf{L}^{\textup{id}}_{\psi_{s_{0}},\psi_{s_{1}}}(\vec{Y}),$
where $\textbf{L}^{\mathcal{G}}_{c_{s_{0}},c_{s_{1}}}(\vec{Y})$ depends on the copula, while $\textbf{L}^{\textup{id}}_{\psi_{s_{0}},\psi_{s_{1}}}(\vec{Y})$, given by (6), is independent of the copula.
Since $\textbf{L}^{\textup{id}}_{\psi_{s_{0}},\psi_{s_{1}}}(\vec{Y})$ is a sum of independent random variables, the privacy loss distribution (PLD) of $\textbf{L}^{\textup{id}}_{\psi_{s_{0}},\psi_{s_{1}}}(\vec{Y})$ corresponds to the convolution of the individual PLDs of $\{\textbf{L}_{\psi_{i}^{s_{0}},\psi_{i}^{s_{1}}}(Y_{i})\}_{i=1}^{k}$, analogous to composition in DP frameworks [21]. In contrast, $\textbf{L}^{\mathcal{G}}_{c_{s_{0}},c_{s_{1}}}(\vec{Y})$, which depends on the CDFs $\{\Psi_{i}\}_{i=1}^{k}$, generally disrupts this property, preventing the PLD of $\textbf{L}_{b_{s_{0}},b_{s_{1}}}(\vec{Y})$ from being a straightforward convolution of the individual PLDs.
Theorem 1 establishes that the composition of individual DCP mechanisms remains DCP but lacks the graceful behavior of DP composition due to the loss of convolution, even for independent mechanisms. The full discussion and quantitative bounds are provided in Appendix -C.
Theorem 1
Let $\mathcal{M}(\vec{\gamma})$ be the composition of $\{\mathcal{M}_{i}(\gamma_{i})\}_{i=1}^{k}$ be independent mechanisms. Let $\mathcal{N}_{i}(\psi_{i})$ be the effective mechanism of $\mathcal{M}_{i}(\gamma_{i})$, $\forall i$. Suppose $\textbf{L}^{\mathcal{G}}_{c_{s_{0}},c_{s_{1}}}(\vec{Y})\neq 0$. Then, the following holds:
  • (i)
    $\mathcal{M}(\vec{\gamma})$ is $(\epsilon_{g},\delta_{g})$-DCP for some $\epsilon_{g}\geq 0$ and $\delta_{g}\in[0,1]$.
  • (ii)
    $\mathcal{M}(\vec{\gamma})$ is not generally $(\sum_{i=1}^{k}\epsilon_{i},\sum_{i=1}^{k}\delta_{i})$-DCP.
  • (iii)
    $\textup{OPT}\left(\mathcal{N}(s;\vec{\psi}),\delta_{g}\right)$ strictly underestimates the privacy risk of $\mathcal{M}(\vec{\gamma})$.
IV Inverse Composition
We propose inverse composition (IC) to address challenges in privacy composition by formulating an optimization framework involving two decision-makers: a leader and a follower. The leader employs a privacy strategy $\alpha:\mathcal{S}\mapsto\Delta(\mathcal{Y}_{\alpha})$, where $\mathcal{Y}_{\alpha}$ is the output space, to achieve any of two tasks:
  1. 1.
    Task-1: Add a new mechanism $\mathcal{M}_{k+1}(\gamma_{\alpha}):\mathcal{X}\mapsto\mathcal{Y}_{\alpha}$ to $\mathcal{M}(\vec{\gamma})$ and ensure the new composition, denoted by $\mathcal{M}(\vec{\gamma},\gamma_{\alpha}):\mathcal{X}\mapsto\vec{\mathcal{Y}}\times\mathcal{Y}_{\alpha}$, satisfies a target $(\epsilon_{g},\delta_{g})$-DCP constraint, where $\gamma_{\alpha}$ is determined by $\alpha$.
  2. 2.
    Task-2: Determine the global privacy parameters $(\epsilon_{g},\delta_{g})$ of the existing composition $\mathcal{M}(\vec{\gamma})=(\mathcal{M}_{1}(\gamma_{1}),\dots,\mathcal{M}_{k}(\gamma_{k}))$ without adding a new mechanism.
By observing $\vec{y}_{\alpha}=(\vec{y},y_{\alpha})\in\vec{\mathcal{Y}}_{\alpha}\equiv\vec{\mathcal{Y}}\times\mathcal{Y}_{\alpha}$, the follower adopts a response strategy $\pi:\vec{\mathcal{Y}}_{\alpha}\mapsto\Delta(\mathcal{S})$, reacting optimally to the leader’s decisions by obtaining a probability distribution (inference) over the secrets. In both tasks, the existing $k$ mechanisms may exhibit arbitrary dependencies, captured by $\mathcal{C}$. Additionally, the mechanism $\mathcal{M}_{k+1}(\gamma_{\alpha})$ can either intrinsically depend on the existing $k$ mechanisms or introduce further dependencies through $\alpha$, e.g., via copula perturbation (see Appendix F). In this section, we focus on the case when $\gamma_{\alpha}$ s independent of $\vec{\gamma}$.
The idea of IC is as follows. The privacy strategy $\alpha$ introduces an additional PLRV $\textbf{L}_{\alpha_{s_{0}},\alpha_{s_{1}}}$, resulting in the total PLRV:
$\textbf{L}^{\alpha}_{s_{0},s_{1}}(\vec{Y},Y_{\alpha})=\textbf{L}_{\hat{b}_{s_{0}},\hat{b}_{s_{1}}}(\vec{Y})+\textbf{L}_{\alpha_{s_{0}},\alpha_{s_{1}}}(Y_{\alpha}),$ (7)
where $\textbf{L}_{\hat{b}_{s_{0}},\hat{b}_{s_{1}}}(\vec{Y})$, given by (12) of Appendix -C, is fixed but unknown. Thus, changes in $\textbf{L}^{\alpha}{s_{0},s_{1}}$ are driven by $\textbf{L}_{\alpha_{s_{0}},\alpha_{s_{1}}}$, with $\alpha$ also influencing the response strategy $\pi$. Traditional forward approaches rely on generally challenging worst-case proofs, including sensitivity calculations (NP-hard [8]) required for the proper distributions of $\alpha$, and combinatorial analysis to compute $\epsilon_{g}$ and $\delta_{g}$ (#P-complete even for independent composition [6]). In contrast, IC designs $\alpha$ to elicit a follower response $\pi$ with certain properties. The key insight is that if $\alpha$ induces a $\pi$ satisfying these properties, the corresponding $\alpha$ will lead to $\textbf{L}^{\alpha}_{s_{0},s_{1}}$ that automatically achieves the desired privacy guarantees, without requiring explicit constraints on $\alpha$.
By doing so, IC ensures that the resulting privacy strategy $\alpha$ inherently fulfills Task-1 and Task-2 without requiring explicit constraints on $\alpha$, thereby avoiding the need for worst-case proofs. We formulate IC as a constrained optimization problem. The leader and the following choose $\alpha$ and $\pi$ by minimizing an expected loss based on any strictly proper scoring rule (SPSR) $\mathcal{L}$ taking the form [30]
$\mathcal{L}(\pi,s,\vec{y}_{\alpha})\equiv\Phi(\pi)+\Upsilon(\pi(s|\vec{y}_{\alpha})),$ (8)
where $\Phi(\cdot)$ is strict convex, and $\Upsilon$ is some pointwise function. In addition, for any valid $p:\vec{\mathcal{Y}}_{\alpha}\to\Delta(\mathcal{S})$, the expected loss satisfies
$\mathbb{E}_{\alpha,\vec{\gamma}}^{p}[\mathcal{L}(p,S,\vec{Y}_{\alpha})]\leq\mathbb{E}_{\alpha,\vec{\gamma}}^{p}[\mathcal{L}(\pi,S,\vec{Y}_{\alpha})],$
where the expectation captures randomness in $S$ and $\vec{Y}_{\alpha}$. It is well-known that $\pi^{*}=p$ is the unique minimizer of $\mathbb{E}_{\alpha,\vec{\gamma}}^{p}[\mathcal{L}(\pi,S,\vec{Y}_{\alpha})]$ [31].
For $\tau_{g}\geq 1$ and $\delta_{g}\in(0,1]$, define:
$\Pi[\tau_{g},\delta_{g}]\equiv\left\{\pi\middle|\begin{aligned} &\pi(s\mid\vec{y}_{\alpha})\geq\tau^{-1}_{g}P_{\theta}(s),\quad\forall s,\vec{y}_{\alpha},\\ &\mathbb{E}_{\pi}\left[\frac{\pi(S\mid\vec{y}_{\alpha})}{P_{\theta}(S)}\right]\leq\delta_{g}\tau_{g},\quad\forall\vec{y}_{\alpha}\end{aligned}\right\},$
where the expectation captures randomness in $S$ by $\pi(\cdot\mid\vec{y}_{\alpha})$. Additionally, for $\delta_{g}=0$, define:
$\Pi[\tau_{g},0]\equiv\left\{\pi\middle|\tau^{-1}_{g}P_{\theta}(s)\leq\pi(s\mid\vec{y}_{\alpha})\leq\tau_{g}P_{\theta}(s),\quad\forall s,\vec{y}_{\alpha}\right\}.$
Convex Optimization Given $\tau_{g}\geq 1$ and $\delta_{g}\in[0,1]$, consider the following two optimization problems:
$\min_{\alpha,\pi}\mathbb{E}^{\mu}_{\alpha,\vec{\gamma}}\left[\mathcal{L}(\pi,S,\vec{Y}_{\alpha})\right],\text{ s.t. }\pi\in\Pi[\tau_{g},\delta_{g}],$ (Task-1)
$\min_{\tau_{g},\alpha,\pi}\mathbb{E}^{\mu}_{\alpha,\vec{\gamma}}\left[\mathcal{L}(\pi,S,\vec{Y}_{\alpha})\right],\text{ s.t. }\pi\in\Pi[\tau_{g},\delta_{g}].$ (Task-2)
The posterior $\mu$ need not be convex in $\alpha$. However, if we choose an SPSR ^$\hat{\mathcal{L}}(\pi,s,\vec{y}_{\alpha})=-\log\bigl{(}\pi(s\mid\vec{y}_{\alpha})\bigr{)}\,P_{\alpha,\vec{\gamma}}(\vec{y}_{\alpha})$, where $P_{\alpha,\vec{\gamma}}(\vec{y}_{\alpha})=\sum_{s}\alpha(y_{\alpha}\mid s)\,P_{\theta}(s)$, then the factor $P_{\alpha,\vec{\gamma}}(\vec{y}_{\alpha})$ cancels inside the expectation $\mathbb{E}^{\mu}_{\alpha,\vec{\gamma}}[\mathcal{L}]$, leaving only the product $\alpha(y_{\alpha}\mid s)\,P_{\theta}(s)$. Moreover, $\Pi[\tau_{g},\delta_{g}]$ is convex in $\pi$ and $\tau_{g}$ and does not depend on $\alpha$. Hence, with $\mathcal{L}=\hat{\mathcal{L}}$, both (Task-1) and (Task-2) reduce to convex programs.
Theorem 2
Given $\{\theta,\mathcal{G}\}$, let $\{\mathcal{M}_{i}(\gamma_{i})\}_{i=1}^{k}$ be $k$ mechanisms, each satisfying $(\epsilon_{i},\delta_{i})$-DCP. Let $\varepsilon(\tau_{g},\theta)\equiv\log\left(1+\frac{\tau_{g}-1}{P^{*}_{\theta}}\right)$ with $P^{*}_{\theta}=\min_{s}P_{\theta}(s)>0$. When $\delta_{g}\neq 0$, let $\alpha^{*}$ be a solution to (Task-1) and $\tau^{*}_{g}$ a solution to (Task-2). Then, for any $\delta_{g}\in[0,1]$, the following hold:
  • (i)
    (Task-1) For any $\tau_{g}\geq 1$, if $\Pi[\tau_{g},\delta_{g}]\neq\emptyset$, the composition $\mathcal{M}(\vec{\gamma},\gamma_{\alpha^{*}})$ is $(\varepsilon(\tau_{g},\theta),\delta_{g})$-DCP.
  • (ii)
    (Task-2) The composition $\mathcal{M}(\vec{\gamma})$ is $(\varepsilon(\tau^{*}_{g},\theta),\delta_{g})$-DCP.
  • (iii)
    When $\delta_{g}=0$, $\mathcal{M}(\vec{\gamma},\gamma_{\alpha^{\dagger}})$ is $\varepsilon(\tau_{g},\theta)$-DCP if $\Pi[\tau_{g},0]\neq\emptyset$, for $\tau_{g}\geq 1$, and $\alpha^{\dagger}$ is a solution to (Task-1). Furthermore, $\mathcal{M}(\vec{\gamma})$ is $\varepsilon(\tau^{\dagger}_{g},\theta)$-DCP if and only if $\tau^{\dagger}_{g}$ is a solution to (Task-2).
Theorem 2 establishes that the privacy strategy $\alpha$ can be designed via IC to achieve Task-1 and Task-2 by eliciting an optimal response strategy $\pi$. When $\delta_{g}\neq 0$, the optimal $\pi$ satisfying the conditions of $\Pi[\tau_{g},\delta_{g}]$ is sufficient for the corresponding $\alpha$ to achieve $(\varepsilon(\tau_{g},\theta),\delta_{g})$-DCP. In Task-1, $\textbf{L}_{\alpha_{s_{0}},\alpha_{s_{1}}}$ is constructed to ensure that $\textbf{L}^{\alpha}_{s_{0},s_{1}}$ satisfies $(\varepsilon(\tau_{g},\theta),\delta_{g})$-DCP without wasting privacy budget. In Task-2, $\textbf{L}_{\alpha_{s_{0}},\alpha_{s_{1}}}=0$, and the optimal $\epsilon^{*}_{g}=\varepsilon(\tau^{*}_{g},\theta)$ represents the tightest privacy parameter for $\textbf{L}^{\alpha}_{s_{0},s_{1}}$. When $\delta_{g}=0$, the optimal $\pi$ satisfying the conditions of $\Pi[\tau_{g},0]$ is sufficient for $\alpha$ to achieve $(\varepsilon(\tau_{g},\theta),0)$-DCP. However, these conclusions assume an ideal scenario where the convex optimization problems can be solved exactly. Since both (Task-1) and (Task-2) are infinite-dimensional, they cannot be solved directly. The exploration of algorithmic and computational methods to address these challenges is left as future work.
IV-A Numerical Experiments
(a) Independent Perturbation
(b) Copula Perturbation
Figure 1: (a) shows five tests with $\delta_{g}=\delta_{i}=0.02$ and $\epsilon_{g}=\{0.25,0.25,1.5,3,5\}$, corresponding to $\epsilon_{i}=\{0.05,0.1,0.3,0.6,1\}$ for $i\in\{1,2,3,4,5\}$. (b) shows six tests with $\delta_{g}=\delta_{i}=0.02$ and $\epsilon_{g}=\{0.4,0.6,1,2,4,6\}$, corresponding to $\epsilon_{i}=\{0.05,0.1,0.18,0.3,0.6,1\}$ for $i\in\{1,2,3,4,5\}$.
In this section, we validate our IC approach through numerical experiments. The constrained optimization problem (inverse composition) is solved using the penalty method. The dataset used is from the 2016 iDASH Workshop on Privacy and Security [32], derived from the 1000 Genomes Project [33], consisting of SNVs from 800 individuals on Chromosome 10. All experiments were conducted on an NVIDIA A40 48G GPU using PyTorch. Detailed experimental settings are provided in Appendix M.
In the first experiment, we demonstrate how IC optimally adds a fifth mechanism, randomized by $\alpha$ using output perturbation, so that the composition of five DCP mechanisms achieves the target privacy budget $(\epsilon_{g},\delta_{g})$. In the second experiment, we show how IC performs copula perturbation (see Appendix A-B) to achieve the target privacy budget $(\epsilon_{g},\delta_{g})$ for five DCP mechanisms. Copula perturbation using $\alpha$: (1) manipulates the dependencies between two chosen mechanisms (e.g., $\mathcal{M}_{1}(\gamma_{1})$ and $\mathcal{M}_{2}(\gamma_{2})$), and (2) transforms $\textbf{L}_{\alpha_{s_{0}},\alpha_{s_{1}}}$ in (7) into a copula PLRV with $\gamma_{1}$ and $\gamma_{2}$ as marginals.
We refer to these experiments as independent perturbation and copula perturbation, respectively. The validity of IC is assessed by comparing the privacy loss of the composed mechanisms to that of a single $(\epsilon_{g},\delta_{g})$-DCP mechanism under the same membership inference attack. Privacy loss is evaluated through the Area Under the Curve (AUC) of membership inference attacks, which audit the extent of data leakage in machine learning models [34, 35]. As we can see from Fig. 1, the AUC under IC-based composition is comparable to that of the single mechanism, confirming that IC efficiently implements the composition.
V Conclusion
We introduced Differential Confounding Privacy (DCP), extending differential privacy to settings involving complex relationships between secrets and datasets. While DCP mechanisms are composable, they lack the graceful composition properties of DP. To address this, we proposed the Inverse Composition (IC) framework, which enables systematic design of privacy strategies. Future work includes algorithmic development of IC and exploration of its approximations.
Appendix A CP As Single-Point DP
Consider that the prior knowledge $\theta\in\Theta$ is a uniform distribution (or other non-informative prior) over $\mathcal{S}$. In addition, let $\mathtt{D}(s,s^{\prime}|\theta)$ represent Hamming distance and $\mathtt{d}=1$. Thus, any adjacent pair $s$ and $s^{\prime}$ with $\mathtt{D}(s,s^{\prime}|\theta)\leq 1$ differ in only one entry (denoted by $s\sim s^{\prime}$). Thus, in this case, the (hypothetical) sensitive information is an entry of a secret, where the relationship between the sensitive information and the secret is deterministic. Then, $(\epsilon,\delta)$-DCP is a notion of $(\epsilon,\delta)$-DP when we treat the dataset $x$ as some intermediate terms in the probabilistic pipeline from the secret to the output.
When we consider adjacent secrets with priors and $\mathcal{G}$ is deterministic and invertible, the DCP framework also mathematically coincides with the Bayesian differential privacy (BDP).
Definition 3 ($(\epsilon_{\omega},\delta_{\omega})$-Bayesian Differential Privacy [37])
A randomized mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}$ is $(\epsilon_{\omega},\delta_{\omega})$-Bayesian differentially private ($(\epsilon_{\omega},\epsilon_{\omega})$-BDP) with $\epsilon_{\omega}\geq 0$ and $\delta_{\omega}\in[0,1]$, if for any $x\simeq x^{\prime}$ differing in a single entry $z\sim\omega$, we have
$\textup{Pr}\left[\mathcal{M}(x)\in\widehat{\mathcal{Y}}\right]\leq e^{\epsilon_{\omega}}\textup{Pr}\left[\mathcal{M}(x^{\prime})\in\widehat{\mathcal{Y}}\right]+\delta_{\omega},\forall\widehat{\mathcal{Y}}\subseteq\mathcal{Y},$
where the probability is taken over the randomness of the output response $y$ and the data entry $z$.
The notion of BDP enjoys many properties of the standard DP [37], including post-processing, composition, and group privacy. The core idea of BDP lies in defining typical scenarios. A scenario is considered as typical when all sensitive data is drawn from the same distribution [37]. This is motivated by the fact that many machine learning models are often designed and trained for specific data distributions, and such prior knowledge is usually known by the attackers. The BDP framework adjusts the noise according to the data distribution and can provide a better expected privacy guarantee.
In the general DCP framework, we treat the entire secret as sensitive information and consider the probability space over all states, while the BDP only considers the space over a single differing entry. Since the common data entries of the adjacent datasets are assumed to be known by the (worst-case) attacker, these common entries do not impact the distributions of the mechanism. In addition, both standard DP and BDP consider independent data.
That is, the privacy guarantees by DP and BDP do not consider the correlation between data entries. Therefore, we can mathematically treat the DCP framework as the BDP framework when the virtual dataset (that contains sensitive entry) is a single-point dataset and the true dataset $x$ is some intermediate term that is not publicly observable. However, such a single-point (B)DP equivalence is valid only when there is one mechanism. When multiple DCP mechanisms are composed, this equivalence fails (See Section III for more detail).
A-A Well-Defined Dependence Structure
In this section, we describe what the well-defined dependence structure is (with slight abuse of notations for simplicity).
D.1 Probability Space and Random Mechanisms
Let $(\mathcal{X},\mathcal{F},\mathbb{P})$ be a probability space, where:
  • $\mathcal{X}$ is the set of possible input datasets,
  • $\mathcal{F}$ is a $\sigma$-algebra of measurable subsets of $\mathcal{X}$,
  • $\mathbb{P}$ is the probability measure defined on $(\mathcal{X},\mathcal{F})$.
Let $M_{1},M_{2},\dots,M_{k}$ be mechanisms (random variables) that introduce randomness to ensure privacy. Each mechanism $M_{i}$ maps a dataset $x\in\mathcal{X}$ and a random input $\omega\in\Omega^{\prime}$ (from a probability space $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime})$) to an output in a measurable space $(\mathcal{Y}_{i},\mathcal{B}_{i})$. Formally:
$M_{i}:\mathcal{X}\times\Omega^{\prime}\to\mathcal{Y}_{i},$
where $\mathcal{B}_{i}\subseteq 2^{\mathcal{Y}_{i}}$ is a $\sigma$-algebra on $\mathcal{Y}_{i}$.
For a fixed dataset $x\in\mathcal{X}$, each $M_{i}(x)$ is a random variable defined as:
$M_{i}(x)=M_{i}(x;\omega),\quad\omega\sim\mathbb{P}^{\prime}.$
  1. 1.
    For a fixed $x\in\mathcal{X}$, denote by $F_{i}^{(x)}$ the (marginal) distribution function (CDF) of $M_{i}(x)$:
    • If $M_{i}(x)$ has a density $f_{i}^{(x)}$, then:
[…]
  2. 2.
    Let $F^{(x)}$ be the joint distribution function of the vector:
    $\mathbf{M}(x)=\bigl{(}M_{1}(x),M_{2}(x),\dots,M_{k}(x)\bigr{)}.$
    The joint CDF $F^{(x)}$ is defined over the product space $(\mathcal{Y}_{1}\times\mathcal{Y}_{2}\times\cdots\times\mathcal{Y}_{k},\mathcal{B}_{1}\otimes\mathcal{B}_{2}\otimes\cdots\otimes\mathcal{B}_{k})$ as:
    $F^{(x)}(m_{1},m_{2},\dots,m_{k})$
    $=\mathbb{P}\big{(}M_{1}(x)\leq m_{1},\dots,M_{k}(x)\leq m_{k}\big{)}.$
A fundamental question in dependence in probability theory and statistics is how and in what ways $F^{x}$ deviates from independence. More generally, how the joint law $F^{x}$ “couples" the marginals $\{F^{x}_{i}\}^{k}_{i=1}$.
D.2 Dependence Structure via Copulas
Probability-Integral Transforms We consider that each marginal $F^{x}_{i}$ is continuous for all $x\in\mathcal{X}$. Define the probability-integral transform for each mechanism $M_{i}(s)$:
$U_{i}(s,\omega)\;:=\;F_{i}^{(s)}\bigl{(}M_{i}(s,\omega)\bigr{)}.$
Then each $U_{i}$ is a random variable taking values in $[0,1]$. Indeed, by the usual probability-integral-transform result, $U_{i}\sim\text{Uniform}(0,1)$ when $M_{i}(s,\omega)$ is continuous.
The Copula as the Joint Law of $\mathbf{U}(s)$ Consider the random vector
$\mathbf{U}(s)$ $=\bigl{(}U_{1}(s),U_{2}(s),\dots,U_{k}(s)\bigr{)}$
$=\bigl{(}F_{1}^{(s)}(M_{1}(s)),\;F_{2}^{(s)}(M_{2}(s)),\;\dots,\;F_{k}^{(s)}(M_{k}(s))\bigr{)}.$
This vector $\mathbf{U}(s)$ lives in the unit cube $[0,1]^{k}$. Define the copula $C^{(s)}$ for the dataset $s\in\mathcal{S}$ as: for $\mathbf{u}\in[0,1]^{k}$,
$C^{(s)}(\mathbf{u})\;:=\;\mathbb{P}\!\Bigl{(}U_{1}(s)\leq u_{1},\,U_{2}(s)\leq u_{2},\,\dots,\,U_{k}(s)\leq u_{k}\Bigr{)}.$
Then $C^{(s)}$ is itself a distribution function on $[0,1]^{k}$, called the copula associated with the dataset $s$.
Sklar’s Theorem By Sklar’s theorem, the original joint distribution $F^{(s)}$ of $\mathbf{M}(s)$ and the marginals $F_{i}^{(s)}$ satisfy:
$F^{(s)}(m_{1},\dots,m_{k})\;=\;C^{(s)}\bigl{(}F_{1}^{(s)}(m_{1}),\dots,F_{k}^{(s)}(m_{k})\bigr{)}.$
Moreover, we have the following.
  • Existence: $C^{(s)}$ always exists for any valid joint $F^{(s)}$ and marginals $F_{i}^{(s)}$.
  • Uniqueness: If each $F_{i}^{(s)}$ is continuous, $C^{(s)}$ is unique. (If some marginals are not continuous, the copula need not be unique but it still exists.)
Thus, from a rigorous standpoint, the dependence structure among $\mathbf{M}(s)$ is precisely the copula $C^{(s)}$—that is, the joint distribution of the transformed vector $\mathbf{U}(s)$.
If some marginal $F_{i}^{(s)}$ is not continuous, we still define
$U_{i}(s)=F_{i}^{(s)}(M_{i}(s)).$
Then each $U_{i}(s)$ is not necessarily $\text{Uniform}(0,1)$ in the usual sense (it may have jumps/atoms), but $\mathbf{U}(s)$ still lives in $[0,1]^{k}$. A copula $C^{(s)}$ satisfying
$F^{(s)}(m_{1},\dots,m_{k})\;=\;C^{(s)}\bigl{(}F_{1}^{(s)}(m_{1}),\dots,F_{k}^{(s)}(m_{k})\bigr{)}$
still exists (by Sklar’s theorem), although it may not be unique when there are atoms.
Putting it all together, a concise definition of well-defined dependence structure is as follows.
Definition 4
Let $\mathbf{M}(s)$ be a $k$-dimensional random vector with continuous marginals $F_{i}^{(s)}$. Define
$U_{i}(s)=F_{i}^{(s)}(M_{i}(s)),$
and let $\mathbf{U}(s)=\bigl{(}U_{1}(s),\dots,U_{k}(s)\bigr{)}$. The joint distribution of $\mathbf{U}(s)$, which is a probability measure on $[0,1]^{k}$, is called the copula $C^{(s)}$. This measure $C^{(s)}$ completely characterizes the dependence structure among the components of $\mathbf{M}(s)$.
Intuitively, the marginals $F_{i}^{(s)}$ “remove” the specific scale and shape of each mechanism’s distribution (making each $U_{i}(s)$ uniform on $[0,1]$). Then, what remains (i.e., how the $U_{i}(s)$ jointly vary in the unit cube) is the pure dependence structure among the mechanisms, independent of the marginal properties.
To summarize, the independence of $\mathbf{M}(s)$ corresponds to the case where the copula $C^{(s)}$ is the product copula, i.e., $C^{(s)}(\mathbf{u})=u_{1}\cdot u_{2}\cdots u_{k}$. Any other copula $C^{(s)}$ indicates non-trivial dependence. If $\mathbf{M}(s)$ has a well-defined joint distribution $F^{(s)}$ with marginals $F_{i}^{(s)}$, a corresponding copula $C^{(s)}$ always exists and uniquely captures the dependence structure (assuming continuous marginals). Specifically, the dependence among $\mathbf{M}(s)$ is represented by the equivalence class of all joint distributions on $(\mathcal{Y}_{1}\times\cdots\times\mathcal{Y}_{k})$ that share the same marginals $F_{i}^{(s)}$. Concretely, this dependence structure is identified with the copula $C^{(s)}$, which is the joint distribution of the transformed vector $\bigl{(}F_{1}^{(s)}(M_{1}(s)),\dots,F_{k}^{(s)}(M_{k}(s))\bigr{)}$.
A-B Gaussian Copula Perturbation
For any $\theta\in\Theta$, we let $\mathcal{Q}\equiv\{(s,s^{\prime})|\mathtt{D}(s,s^{\prime}|\theta)\leq\mathtt{d}\}$ for simplicity. We start by defining bivariate Gaussian copula.
Definition 5 (Bivariate Gaussian Copula)
A bivariate copula $C^{\mathtt{g}}_{\rho}:[0,1]^{2}\to[0,1]$ is called a Gaussian copula if it is defined as
$C^{\mathtt{g}}_{\rho}(u_{1},u_{2})=\Phi_{\rho}\big{(}\Phi^{-1}(u_{1}),\Phi^{-1}(u_{2})\big{)},$
where:
  • $\Phi$ is the CDF of the standard normal distribution, mapping $u\in[0,1]$ to $\mathbb{R}$ via its inverse $\Phi^{-1}$,
  • $\Phi_{\rho}$ is the joint CDF of a bivariate normal distribution with zero means, unit variances, and correlation coefficient $\rho$.
This construction ensures $C^{\mathtt{g}}_{\rho}$ satisfies the properties of a copula, mapping $[0,1]^{2}\to[0,1]$ while capturing dependence structure via $\rho$.
Pseudo-random Sample Generation
Let $\xi_{1}$ and $\xi_{2}$ be the density functions of two independent noise distributions $V_{1}\in\mathcal{V}_{1}$ and $V_{2}\in\mathcal{V}_{2}$, with corresponding CDFs $\Xi_{1}$ and $\Xi_{2}$. Additionally, let $Z_{1}\sim N(\mu_{1},\mathtt{var}_{1})$ and $Z_{2}\sim N(0,1)$, with $\textup{F}_{1}(\cdot)$ as the CDF of $N(\mu_{1},\mathtt{var}_{1})$. For any $\rho\neq 0$, the pseudo-random sample generation process is as follows:
  1. 1.
    Sample Generation: Generate $(z_{1},z_{2})$ where $z_{1}\sim N(\mu_{1},\mathtt{var}_{1})$ and $z_{2}\sim N(0,1)$.
  2. 2.
    Transform to Uniform Samples: Compute $u_{1}=\textup{F}_{1}(z_{1})$ and
    $u_{2}=\textup{F}_{\rho}\big{(}\rho z_{1}+\sqrt{1-\rho^{2}}z_{2}\big{)},$
    where $\textup{F}_{\rho}$ is the CDF of $N\big{(}\rho\mu_{1},\rho^{2}\mathtt{var}_{1}+(1-\rho^{2})\big{)}$.
  3. 3.
    Transform to Noise Samples: Obtain $v_{1}=\Xi^{-1}_{1}(u_{1})$ and $v_{2}=\Xi^{-1}_{2}(u_{2})$, where $\Xi^{-1}_{1}$ and $\Xi^{-1}_{2}$ are the inverses of the CDFs corresponding to $\xi_{1}$ and $\xi_{2}$, respectively.
Let $\textup{PsedR}_{\rho}(\cdot|\Xi_{1},\Xi_{2}):\mathbb{R}\times\mathbb{R}\to\mathcal{V}_{1}\times\mathcal{V}_{2}$ denote the deterministic mapping from $(z_{1},z_{2})$ to $(v_{1},v_{2})$ via the steps outlined above, such that
$(v_{1},v_{2})=\textup{PsedR}_{\rho}(z_{1},z_{2}|\Xi_{1},\Xi_{2}).$
Additionally, let $\overline{\textup{PR}}_{\rho}(Z_{1},Z_{2}|\Xi_{1},\Xi_{2})$ represent the joint distribution of $V_{1}$ and $V_{2}$ induced by $N(\mu_{1},\mathtt{var}_{1})$ and $N(0,1)$ through this process.
Copula Perturbation
Following the pseudo-random sample generation, the distribution $\overline{\textup{PR}}_{\rho}(Z_{1},Z_{2}|\Xi_{1},\Xi_{2})$ has a CDF given by the copula:
$C^{\mathtt{g}}_{\rho}(V_{1},V_{2}|\Xi_{1},\Xi_{2})=\Phi_{\rho}\left(F_{1}^{-1}(\Xi_{1}(V_{1})),F_{\rho}^{-1}(\Xi_{2}(V_{2}))\right),$
where $\Phi_{\rho}$ is the CDF of a bivariate normal distribution with correlation $\rho$.
We say that a pair of mechanisms $\mathcal{M}_{1}(\gamma_{1})$ and $\mathcal{M}_{2}(\gamma_{2})$ are perturbed by the copula $C^{\mathtt{g}}_{\rho}(\Xi_{1},\Xi_{2})$ if the noises $v_{1}$ and $v_{2}$ used to perturb $\mathcal{M}_{1}(\gamma_{1})$ and $\mathcal{M}_{2}(\gamma_{2})$, respectively, are sampled via
$(v_{1},v_{2})=\textup{PsedR}_{\rho}(z_{1},z_{2}|\Xi_{1},\Xi_{2}).$
Consider an arbitrary function $\tilde{\eta}_{S}:\mathcal{S}\to\mathcal{B}$ that produces a univariate output $\tilde{\eta}_{S}(s)$ for a given state $s$. Let the sensitivity of $\tilde{\eta}_{S}$ be defined as $\mathtt{C}_{\textup{sen}}=\sup_{(s_{1},s_{0})\in\mathcal{Q}}|\tilde{\eta}_{S}(s_{0})-\tilde{\eta}_{S}(s_{1})|$.
Suppose that for any $s\in\mathcal{S}$,
$z_{1}(s)=\tilde{\eta}_{S}(s)+\tilde{z}_{1},$
where $\tilde{z}_{1}$ is drawn from $N\left(0,\mathtt{w}^{2}\left(\frac{\mathtt{C}_{\textup{sen}}}{\epsilon_{\mathtt{c}}}\right)^{2}\right)$ with $\mathtt{w}\geq 2\log(2/\delta_{\mathtt{c}})$, $\epsilon_{\mathtt{c}}\geq 0$, and $\delta_{\mathtt{c}}\in[0,1)$, and $\tilde{z}_{2}$ is drawn from $N(0,1)$.
Let $\Xi_{1}$ and $\Xi_{2}$ be the CDFs of two independent noise distributions inducing $\gamma_{1}$ and $\gamma_{2}$, respectively, such that the mechanisms $\mathcal{M}_{1}(\gamma_{1})$ and $\mathcal{M}_{2}(\gamma_{2})$ are $(\epsilon_{1},\delta_{1})$-CP and $(\epsilon_{2},\delta_{2})$-CP, respectively. Finally, let $C^{\mathtt{g}}_{\rho}(\Xi_{1},\Xi_{2};\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})$ denote the induced copula.
We then obtain the following theorem.
Theorem 4
Suppose $\mathcal{M}_{1}(\gamma_{1})$ and $\mathcal{M}_{2}(\gamma_{2})$ are perturbed by the copula $C^{\mathtt{g}}_{\rho}(\Xi_{1},\Xi_{2};\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})$ for some $\epsilon_{\mathtt{c}}\geq 0$ and $\delta_{\mathtt{c}}\in[0,1)$. Then, the following holds:
  • (i)
    $\mathcal{M}_{1}(\gamma_{1})$ and $\mathcal{M}_{2}(\gamma_{2})$ remain $(\epsilon_{1},\delta_{1})$-CP and $(\epsilon_{2},\delta_{2})$-CP, respectively.
  • (ii)
    For the composition of $\mathcal{M}_{1}(\gamma_{1})$ and $\mathcal{M}_{2}(\gamma_{2})$,
    $\overline{\textbf{opt}}_{\delta_{g}}=\textup{OptComp}((\epsilon_{\mathtt{c}},\delta_{\mathtt{c}}),\epsilon_{1},\delta_{1},\epsilon_{2},\delta_{2},\delta_{g}),$
    for any $(s_{1},s_{0})\in\mathcal{Q}$ and $\delta_{g}\in[0,1)$, where OptComp is given by Theorem 1.5 of [6].
Consider a collection of independent mechanisms $\{\mathcal{M}_{i}(\gamma_{i})\}_{i}$, where each $\mathcal{M}_{i}(\gamma_{i})$ is $(\epsilon_{i},\delta_{i})$-CP, guaranteed by noise perturbation with distribution $\Xi_{i}$.
Definition 6 (Invertible Noise Perturbation)
A noise perturbation scheme is invertible if there exists a bijective mapping $\mathtt{A}:\mathcal{V}_{i}\to\mathcal{Y}_{i}$ such that $y_{i}=\mathtt{A}(v_{i})$ if and only if $v_{i}=\mathtt{A}^{-1}(y_{i})$, for $i\in\{1,2\}$.
Without loss of generality, select $\mathcal{M}_{1}(\gamma_{1})$ and $\mathcal{M}_{2}(\gamma_{2})$ such that they are perturbed by the copula $C^{\mathtt{g}}_{\rho}(\Xi_{1},\Xi_{2};\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})$. Additionally, define
$\textbf{L}_{s_{0},s_{1}}(Y_{1},Y_{2};c^{\mathtt{g}}_{\rho})\equiv\log\frac{c^{\mathtt{g}}_{\rho}(\Psi_{1}(Y_{1}\mid s_{1}),\Psi_{2}(Y_{2}\mid s_{1})\mid\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})}{c^{\mathtt{g}}_{\rho}(\Psi_{1}(Y_{1}\mid s_{0}),\Psi_{2}(Y_{2}\mid s_{0})\mid\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})},$
where $c^{\mathtt{g}}_{\rho}(\cdot\mid\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})$ is the Gaussian copula density of $C^{\mathtt{g}}_{\rho}$, and $(Y_{1},Y_{2})\sim c^{\mathtt{g}}_{\rho}(\Psi_{1}(Y_{1}\mid s_{1}),\Psi_{2}(Y_{2}\mid s_{1}))$.
Proposition 3
Suppose the noise perturbation is invertible. Under the composition of $k$ mechanisms $\{\mathcal{M}_{j}(\gamma_{j})\}_{j=1}^{k}$, where $\{\mathcal{M}_{1}(\gamma_{1}),\mathcal{M}_{2}(\gamma_{2})\}$ are perturbed by the copula $C^{\mathtt{g}}_{\rho}(\Xi_{1},\Xi_{2};\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})$, the privacy loss random variable (PLRV) is
$\widetilde{\mathbf{L}}_{b_{s_{0}},b_{s_{1}}}(\vec{Y};C^{\mathtt{g}}_{\rho})=\textbf{L}_{b_{s_{0}},b_{s_{1}}}(\vec{Y})+\textbf{L}_{s_{0}\|s_{1}}(Y_{1},Y_{2};c^{\mathtt{g}}_{\rho}),$
[…]
Proposition 3 demonstrates that the final PLRV can be additively decomposed into the PLRV induced by the copula perturbation and the unperturbed PLRV of the composition.
In the copula perturbation, we use $N\left(0,\mathtt{w}^{2}\left(\frac{\mathtt{C}_{\textup{sen}}}{\epsilon_{\mathtt{c}}}\right)^{2}\right)$ with $\mathtt{w}\geq 2\log(2/\delta_{\mathtt{c}})$ and $N(0,1)$ to generate the latent variables $z_{1}$ and $z_{2}$, respectively. Additionally, the mechanism $\tilde{\eta}_{S}$ introduces state dependence, where the latent variable $z_{1}$ acts as Gaussian noise to perturb the output of $\tilde{\eta}_{S}$. As a result, this Gaussian copula achieves $(\epsilon_{\mathtt{s}},\delta_{\mathtt{c}})$-CP and contributes to the total cumulative privacy loss of the mechanisms under composition.
The copula perturbation is straightforward to implement, with state dependence captured entirely by the deterministic mechanism $\tilde{\eta}_{S}$, independent of the underlying Gaussian copula distributions. Alternative approaches to copula perturbation are also possible. For example, the parameters of the bivariate Gaussian distribution (e.g., correlation coefficient $\rho$, mean, or variance) used in the copula framework could be selected based on the states, introducing state dependence directly into the copula.
During the pseudo-random sample generation process, noises $v_{1}$ and $v_{2}$ are jointly sampled to perturb the selected mechanisms $\mathcal{M}_{1}(\gamma_{1})$ and $\mathcal{M}_{2}(\gamma_{2})$. However, the copula perturbation does not require these two mechanisms to operate simultaneously. The key requirement is that the noise used to perturb each mechanism is generated via the pseudo-random sample generation process.
A-C Proof of Theorem 4 in Appendix A-B
Part (i) of Theorem 4 demonstrates that our modified Gaussian copula $C^{\mathtt{g}}_{\rho}(\Xi_{1},\Xi_{2};\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})$ is a valid copula that retains the marginals as $\gamma_{1}$ and $\gamma_{2}$.
Part (ii) of Theorem 4 shows that the conservative upper bound provided by Theorem 3 can be computed using Theorem 1.5 of [6], given $(\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})$, which parameterize the Gaussian distribution of $Z_{1}$ in the process $\textup{PsedR}_{\rho}(\cdot\mid\Xi_{1},\Xi_{2})$.
For simplicity, we omit $\epsilon_{\mathtt{c}}$ and $\delta_{\mathtt{c}}$, and write $C^{\mathtt{g}}_{\rho}(\Xi_{1},\Xi_{2})$.
A-C1 Proof of Part (i)
Part (i) of Theorem 4 can be proved by showing that
$C^{\mathtt{g}}_{\rho}(V_{1},V_{2}\mid\Xi_{1},\Xi_{2})=\Phi_{\rho^{\prime}}\left(F_{1}^{-1}(\Xi_{1}(V_{1})),F_{\rho}^{-1}(\Xi_{2}(V_{2}))\right)$
is a valid copula for any $\rho^{\prime}\in(-1,1)$ and $\rho\in(-1,1)\setminus\{0\}$, ensuring that the marginals of the joint distribution $\overline{\textup{PR}}(Z_{1},Z_{2}\mid\Xi_{1},\Xi_{2})$ are $\Xi_{1}$ and $\Xi_{2}$.
Let
$C^{g}_{\rho}(U_{1},U_{2})=\Phi_{\rho^{\prime}}\left(F_{1}^{-1}(U_{1}),F_{\rho}^{-1}(U_{2})\right).$
The function $C^{\mathtt{g}}_{\rho}(V_{1},V_{2}\mid\Xi_{1},\Xi_{2})$ is a valid copula if the following conditions hold:
  • $C^{g}_{\rho}(U_{1},U_{2})$ is a valid joint CDF.
  • $C^{g}_{\rho}(U_{1},U_{2})$ has uniform marginals.
Condition (a) follows directly from the definition of $\Phi_{\rho^{\prime}}$, which is the joint CDF of a bivariate Gaussian distribution with zero mean and correlation coefficient $\rho^{\prime}$.
Uniform Marginal for $U_{1}$ Given any $s$, the CDF of $z_{1}(s)$ is given by
$F_{1}(z_{1}(s))=\Phi\left(\frac{z_{1}-\tilde{\eta}_{S}(s)}{\sqrt{\mathtt{var}_{1}}}\right),$
where $\Phi$ is the CDF of the standard Gaussian distribution, and $\mathtt{var}_{1}=\mathtt{w}^{2}\left(\frac{\mathtt{C}_{\textup{sen}}}{\epsilon_{\mathtt{c}}}\right)^{2}$. By the probability integral transform, $U_{1}=F_{1}(Z_{1}(s))$ is uniformly distributed on $[0,1]$ because $F_{1}$ maps the real line to $[0,1]$ and is a continuous, strictly increasing function.
Uniform Marginal for $U_{2}$ Given any $s$, $F_{\rho}$ is the CDF of the random variable
$\hat{Z}_{2}(s)=\rho Z_{1}(s)+\sqrt{1-\rho^{2}}Z_{2},$
where $Z_{2}\sim N(0,1)$. Therefore,
$F_{\rho}(\hat{z}_{2}(s))=\Phi\left(\frac{\hat{z}_{2}(s)-\rho\tilde{\eta}_{S}(s)}{\sqrt{\rho^{2}\mathtt{var}_{1}+(1-\rho^{2})}}\right).$
By the probability integral transform, $U_{2}=F_{\rho}(\hat{Z}_{2}(s))$ is uniformly distributed on $[0,1]$ because $F_{\rho}$ is a valid CDF that maps $\hat{Z}_{2}(s)$ to a uniformly distributed random variable on $[0,1]$.
Therefore, $C^{\mathtt{g}}_{\rho}(V_{1},V_{2}\mid\Xi_{1},\Xi_{2})$ is a valid copula with marginal distributions $\Xi_{1}$ and $\Xi_{2}$. Here, each $\Xi_{i}$ is independently chosen (without considering composition) such that the mechanism $\mathcal{M}_{i}(\gamma_{i})$ is $(\epsilon_{i},\delta_{i})$-CP for $i\in\{1,2\}$.
Furthermore, if an attacker only has access to one of the two mechanisms, each $\mathcal{M}_{i}(\gamma_{i})$ of the $C^{\mathtt{g}}_{\rho}(V_{1},V_{2}\mid\Xi_{1},\Xi_{2})$-perturbed $\{\mathcal{M}_{1}(\gamma_{1}),\mathcal{M}_{2}(\gamma_{2})\}$ maintains $(\epsilon_{i},\delta_{i})$-CP.
A-C2 Proof of Part (ii)
Recall that $\mathtt{C}_{\textup{sen}}=\sup_{(s_{1},s_{0})\in\mathcal{Q}}|\tilde{\eta}_{S}(s_{0})-\tilde{\eta}_{S}(s_{1})|$ is the sensitivity of a function $\tilde{\eta}_{S}:\mathcal{S}\to\mathcal{B}$, which outputs a univariate value $\tilde{\eta}_{S}(s)$ given the state $s$.
For any state $s\in\mathcal{S}$, let
$z_{1}(s)=\tilde{\eta}_{S}(s)+\tilde{z}_{1},$
where $\tilde{z}_{1}$ is drawn from $N\left(0,\mathtt{w}^{2}\left(\frac{\mathtt{C}_{\textup{sen}}}{\epsilon_{\mathtt{c}}}\right)^{2}\right)$, with $\mathtt{w}\geq 2\log(2/\delta_{\mathtt{c}})$, $\epsilon_{\mathtt{c}}\geq 0$, and $\delta_{\mathtt{c}}\in[0,1)$.
It is straightforward to verify that $z_{1}(s_{0})$ and $z_{1}(s_{1})$ are normally distributed with means $\tilde{\eta}_{S}(s_{0})$ and $\tilde{\eta}_{S}(s_{1})$, respectively, and a common variance
$\mathtt{var}_{1}=\mathtt{w}^{2}\left(\frac{\mathtt{C}_{\textup{sen}}}{\epsilon_{\mathtt{c}}}\right)^{2}.$
For any $s$, let
$C^{g}_{\rho}(U_{1}(s),U_{2}(s))=\Phi_{\rho^{\prime}}\left(F^{-1}_{1}(U_{1}),F^{-1}_{\rho}(U_{2})\right),$
where $U_{1}(s)=F_{1}(Z_{1}(s))$ and $U_{2}(s)=F_{\rho}(\hat{Z}_{2}(s))$.
By Theorem 4.3 in [4], $z_{1}(s)$ satisfies $(\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})$-SDP. By the post-processing property of differential privacy, $C^{g}_{\rho}(U_{1}(s),U_{2}(s))$ is also $(\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})$-SDP.
Here, the formulation $\hat{Z}_{2}(s)=\rho Z_{1}(s)+\sqrt{1-\rho^{2}}Z_{2}$ is deterministic and represents a linear combination of $Z_{1}(s)$ and $Z_{2}$, where $\rho\neq-1$ or $1$ to avoid degenerate cases. The CDFs are strictly monotonic functions. In addition, all inverse CDFs are assumed to be continuous and well-defined.
Therefore, the process $C^{g}_{\rho}(U_{1}(s),U_{2}(s))$ is a deterministic one-to-one mapping from $(z_{1}(s),z_{2})$. Thus, the parameters $(\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})$ of $C^{g}_{\rho}(U_{1}(s),U_{2}(s))$’s CP are tight given $(z_{1}(s),z_{2})$.
Recall that the privacy loss random variable can be defined as
$\widehat{\textbf{L}}_{s_{0},s_{1}}(U_{1}(s),U_{2}(s),Y_{1},Y_{2})$ $\equiv\textbf{L}_{\hat{c}_{s_{0}},\hat{c}_{s_{1}}}(U_{1}(s),U_{2}(s))$
$\quad+\sum_{i=1}^{2}\textbf{L}_{\psi_{i}^{s_{0}},\psi_{i}^{s_{1}}}(Y_{i}),$
where $\hat{c}_{s}=c^{\mathtt{g}}_{\rho}(U_{1}(s),U_{2}(s))$.
If the copula is treated as an independent mechanism accessing any dataset related to $s$, we can derive the conservative bound
$\overline{\textbf{opt}}_{\delta_{g}}=\textup{OptComp}((\epsilon_{\mathtt{c}},\delta_{\mathtt{c}}),\epsilon_{1},\delta_{1},\epsilon_{2},\delta_{2},\delta_{g}),$
for any $(s_{1},s_{0})\in\mathcal{Q}$ and $\delta_{g}\in[0,1)$, where OptComp is provided by Theorem 1.5 of [6].
$\square$
A-D Proof of Proposition 3 in Appendix A-B
For simplicity, we omit $(\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})$ and write $C^{\mathtt{g}}_{\rho}(\Xi_{1},\Xi_{2};\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})$.
Recall that $\psi_{i}:\mathcal{S}\to\Delta(\mathcal{Y}_{i})$ is the density function induced by $\gamma_{i}$ and $\theta$, and $\Psi_{i}$ is the corresponding CDF.
Suppose the noise perturbation is given by $y_{i}=\mathtt{A}(v_{i})$, where $\mathtt{A}$ is strictly increasing. Then, $\mathtt{A}^{-1}(Y_{i})\leq v_{i}$ is equivalent to $Y_{i}\leq\mathtt{A}(v_{i})$. Thus, we have
$\Xi_{i}(v_{i})=\textup{Pr}\left(\mathtt{A}^{-1}(Y_{i})\leq v_{i}\right)=\Psi_{i}(\mathtt{A}(v_{i})\mid s).$
Similarly, suppose the noise perturbation is given by $y_{i}=\mathtt{A}(v_{i})$, where $\mathtt{A}$ is strictly decreasing. Then, $\mathtt{A}^{-1}(Y_{i})\leq v_{i}$ is equivalent to $Y_{i}\geq\mathtt{A}(v_{i})$. Thus, we have
$\Xi_{i}(v_{i})=\textup{Pr}\left(\mathtt{A}^{-1}(Y_{i})\geq v_{i}\right)=1-\Psi_{i}(\mathtt{A}(v_{i})\mid s).$
Therefore, when the invertible $\mathtt{A}$ is strictly increasing or strictly decreasing, we have:
$\widehat{C}^{\mathtt{g}}_{\rho^{\prime}}(V_{1},V_{2}\mid\Xi_{1},\Xi_{2};s)$
$=\Phi_{\rho^{\prime}}\left(F_{1}^{-1}(\Psi_{1}(\mathtt{A}(V_{1})\mid s)),F_{\rho}^{-1}(\Psi_{2}(\mathtt{A}(V_{2})\mid s))\right),$
and
$\widetilde{C}^{\mathtt{g}}_{\rho^{\prime}}(V_{1},V_{2}\mid\Xi_{1},\Xi_{2};s)$
$=\Phi_{\rho^{\prime}}\left(F_{1}^{-1}(1-\Psi_{1}(\mathtt{A}(V_{1})\mid s)),F_{\rho}^{-1}(1-\Psi_{2}(\mathtt{A}(V_{2})\mid s))\right).$
In addition, let $\widehat{c}^{\mathtt{g}}_{\rho^{\prime}}(V_{1},V_{2}\mid s)$ and $\widetilde{c}^{\mathtt{g}}_{\rho^{\prime}}(V_{1},V_{2}\mid s)$ denote the corresponding copula densities.
For the two mechanisms $\mathcal{M}_{1}(\gamma_{1})$ and $\mathcal{M}_{2}(\gamma_{2})$ perturbed by $C^{\mathtt{g}}_{\rho}(V_{1},V_{2}\mid\Xi_{1},\Xi_{2})$, let $\Psi_{[2]}(Y_{1},Y_{2}\mid s,C^{\mathtt{g}}_{\rho})$ denote the joint CDF of the outputs $Y_{1}$ and $Y_{2}$.
Lemma 1
The following holds:
  • (i)
    If the invertible $\mathtt{A}$ is strictly increasing, then
    $\Psi_{[2]}(Y_{1},Y_{2}\mid s,C^{\mathtt{g}}_{\rho})=\widehat{C}^{\mathtt{g}}_{\rho}(Y_{1},Y_{2}\mid\Xi_{1},\Xi_{2};s).$
  • (ii)
    If the invertible $\mathtt{A}$ is strictly decreasing, then
    $\Psi_{[2]}(Y_{1},Y_{2}\mid s,C^{\mathtt{g}}_{\rho})=\widetilde{C}^{\mathtt{g}}_{\rho}(Y_{1},Y_{2}\mid\Xi_{1},\Xi_{2};s).$
Proof:
Suppose the invertible $\mathtt{A}$ is strictly increasing. Then,
$\Psi_{[2]}(y_{1},y_{2}\mid s,C^{\mathtt{g}}_{\rho})$ $=\textup{Pr}\left(Y_{1}\leq y_{1},Y_{2}\leq y_{2}\right)$
$=\textup{Pr}\left(\mathtt{A}(V_{1})\leq y_{1},\mathtt{A}(V_{2})\leq y_{2}\right)$
$=\textup{Pr}\left(V_{1}\leq\mathtt{A}^{-1}(y_{1}),V_{2}\leq\mathtt{A}^{-1}(y_{2})\right).$
Since $V_{i}\leq\mathtt{A}^{-1}(y_{i})$ is equivalent to $Y_{i}=\mathtt{A}(V_{i})\leq y_{i}$, we have
$\Xi_{i}(\mathtt{A}^{-1}(y_{i}))=\textup{Pr}\left(V_{i}\leq\mathtt{A}^{-1}(y_{i})\right)=\Psi_{i}(y_{i}),$
for all $i\in\{1,2\}$.
Since $V_{1}$ and $V_{2}$ have the joint CDF $C^{\mathtt{g}}_{\rho}(\cdot\mid\Xi_{1},\Xi_{2})$, it follows that
$\Psi_{[2]}(y_{1},y_{2}\mid s,C^{\mathtt{g}}_{\rho})=C^{\mathtt{g}}_{\rho}(\mathtt{A}^{-1}(y_{1}),\mathtt{A}^{-1}(y_{2})\mid\Xi_{1},\Xi_{2})$
$=$ $\Phi_{\rho^{\prime}}\left(F_{1}^{-1}(\Psi_{1}(\mathtt{A}(\mathtt{A}^{-1}(y_{1}))\mid s)),F_{\rho}^{-1}(\Psi_{2}(\mathtt{A}(\mathtt{A}^{-1}(y_{2}))\mid s))\right)$
$=$ $\Phi_{\rho^{\prime}}\left(F_{1}^{-1}(\Psi_{1}(y_{1}\mid s)),F_{\rho}^{-1}(\Psi_{2}(y_{2}\mid s))\right).$
A similar procedure can be used to prove (ii). ∎
Lemma 1 shows that the dependency between $Y_{1}$ and $Y_{2}$ due to the copula perturbation can be fully characterized by the same structure $C^{\mathtt{g}}_{\rho^{\prime}}$ as the noises $V_{1}$ and $V_{2}$.
Under the copula perturbation using $C^{\mathtt{g}}_{\rho^{\prime}}$, let $b(\cdot\mid s;c^{\mathtt{g}}_{\rho^{\prime}})$ denote the joint density. Then,
$b\left(\vec{Y}\mid s\right)$
$=c\left(\Psi_{[2]}(Y_{1},Y_{2}\mid s,C^{\mathtt{g}}_{\rho}),\Psi_{3}(Y_{3}\mid s),\dots,\Psi_{k}\left(Y_{k}\mid s\right)\mid s\right)$
×$\quad\times\psi_{[2]}(Y_{1},Y_{2}\mid s,C^{\mathtt{g}}_{\rho})\prod_{i=3}^{k}\psi_{i}\left(Y_{i}\mid s\right)$
$=c\left(\Psi_{[2]}(Y_{1},Y_{2}\mid s,C^{\mathtt{g}}_{\rho}),\Psi_{3}(Y_{3}\mid s),\dots,\Psi_{k}\left(Y_{k}\mid s\right)\mid s\right)$
×$\quad\times c^{\mathtt{g}}_{\rho}(y_{1},y_{2}\mid s)\prod_{i=1}^{k}\psi_{i}\left(Y_{i}\mid s\right),$
where $\psi_{[2]}$ is the joint density of the joint CDF $\Psi_{[2]}$.
For any $(s_{0},s_{1})\in\mathcal{Q}$, define
$\widehat{\textbf{L}}_{s_{0},s_{1}}(Y_{1},Y_{2};\widehat{c}^{\mathtt{g}}_{\rho})\equiv\log\frac{\widehat{c}^{\mathtt{g}}_{\rho}(\Psi_{1}(Y_{1}\mid s_{1}),\Psi_{2}(Y_{2}\mid s_{1})\mid\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})}{\widehat{c}^{\mathtt{g}}_{\rho}(\Psi_{1}(Y_{1}\mid s_{0}),\Psi_{2}(Y_{2}\mid s_{0})\mid\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})},$
and
$\widetilde{\textbf{L}}_{s_{0},s_{1}}(Y_{1},Y_{2};\widetilde{c}^{\mathtt{g}}_{\rho})\equiv\log\frac{\widetilde{c}^{\mathtt{g}}_{\rho}(\Psi_{1}(Y_{1}\mid s_{1}),\Psi_{2}(Y_{2}\mid s_{1})\mid\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})}{\widetilde{c}^{\mathtt{g}}_{\rho}(\Psi_{1}(Y_{1}\mid s_{0}),\Psi_{2}(Y_{2}\mid s_{0})\mid\epsilon_{\mathtt{c}},\delta_{\mathtt{c}})}.$
By the product rule for logarithms, we can separate the copula densities $\widehat{c}^{\mathtt{g}}_{\rho}$ and $\widetilde{c}^{\mathtt{g}}_{\rho}$, obtaining
$\widetilde{\textbf{L}}_{b_{s_{0}},b_{s_{1}}}(\vec{Y};C^{\mathtt{g}}_{\rho})=\textbf{L}_{b_{s_{0}},b_{s_{1}}}(\vec{Y})+\textbf{L}_{s_{0},s_{1}}(Y_{1},Y_{2};\widehat{c}^{\mathtt{g}}_{\rho}),$
and
$\widetilde{\textbf{L}}_{b_{s_{0}},b_{s_{1}}}(\vec{Y};C^{\mathtt{g}}_{\rho})=\textbf{L}_{b_{s_{0}},b_{s_{1}}}(\vec{Y})+\widetilde{\textbf{L}}_{s_{0},s_{1}}(Y_{1},Y_{2};\widetilde{c}^{\mathtt{g}}_{\rho}).$
$\square$
A-E Proof of Proposition 1
When $\mathcal{G}$ is an invertible mapping, for any secret $s\in\mathcal{S}$, let $x_{s}\in\mathcal{X}$ be the dataset such that $P_{\theta}(x_{s}|s)=\frac{P_{\theta}(s,x_{s})}{P_{\theta}(s)}=1$. Then, the density function of the effective mechanism $\mathcal{N}(\psi)$ becomes
[…]
$=\gamma_{i}(y_{i}\mid x_{s}).$
The joint probability under composition given by (10) becomes
$\textbf{Pr}^{s}_{\vec{\gamma}}\left[\vec{y}\in\vec{\mathcal{W}}\right]$ $=\int_{x\in\mathcal{X}}\int_{\vec{y}\in\vec{\mathcal{W}}}\prod_{i=1}^{k}\gamma_{i}(y_{i}\mid x)\frac{P_{\theta}(s,x)}{P_{\theta}(s)}\,d\vec{y}\,dx$
$=\int_{\vec{y}\in\vec{\mathcal{W}}}\prod_{i=1}^{k}\gamma_{i}(y_{i}\mid x_{s})d\vec{y}.$
Thus, the joint density under composition becomes the product of individual densities $\vec{\gamma}$. Then, the PLRV of the composition becomes the sum of individual PLRVs. That is,
$\textbf{L}_{b_{s_{0}},b_{s_{1}}}(\vec{Y})=\textbf{L}^{\textup{id}}_{\psi_{s_{0}},\psi_{s_{1}}}(\vec{Y}),$
which implies $\textbf{L}^{\mathcal{G}}_{c_{s_{0}},c_{s_{1}}}(\vec{Y})=0$.
$\square$
A-F Proof of Part (i) of Theorem 1 and Proposition 2
For any $\theta\in\Theta$, let $\mathcal{Q}\equiv\{(s,s^{\prime})|\mathtt{D}(s,s^{\prime}|\theta)\leq\mathtt{d}\}$ for simplicity. Our proof depends on a hypothesis test result.
A-F1 Hypothesis Testing
The notion of $f$-differential privacy [20] generalizes the differential privacy by interpreting the worst-case scenario differential privacy as a Neyman-Pearson optimal hypothesis test. It is not hard to extend $f$-differential privacy ($f$-DP) to our confounding privacy settings. For $(s_{0},s_{1})\in\mathcal{Q}$, consider the following binary hypothesis testing:
$H_{0}$: the state is $s_{0}$. versus $H_{1}$: the state is $s_{1}$.
We rewrite $\textbf{Pr}^{\theta,s}_{\gamma}\left[y\in\mathcal{W}\right]$ in (2) by $\textbf{Pr}\left[\cdot\middle|H_{t},\gamma\right]$, for $t\in\{0,1\}$. For any rejection rule $\phi$, define the Type-I error $\alpha_{\phi}$ and the Type-II error $\beta_{\phi}$ by
$\alpha_{\phi}=\mathbb{E}[\phi|H_{0},\gamma],\beta_{\phi}=1-\mathbb{E}[\phi|H_{1},\gamma].$
Definition 7 (Optimal Symmetric Trade-off Function [20])
A function $f:[0,1]\mapsto[0,1]$ is a symmetric trade-off function if it is convex, continuous, non-increasing, $f(\alpha)\leq 1-\alpha$, and $f(\alpha)=\inf\{t\in[0,1]:f(t)\leq\alpha\}$ for $\alpha\in[0,1]$.
Lemma 2
Fix any $\alpha\in[0,1]$. For any $\theta\in\Theta$, there exists a symmetric trade-off function $f$ and $\epsilon\geq 0$ for the mechanism $\mathcal{M}(\gamma)$ such that $\mathcal{M}$ is $(\epsilon,\delta(\epsilon))$-DCP, where
$\delta(\epsilon)=1+f^{*}(-e^{\epsilon})\textup{ with }f^{*}(y)=\sup\nolimits_{0\leq\alpha^{\prime}\leq 1}y\alpha^{\prime}-f(\alpha^{\prime}).$
Proof:
Lemma 2 is a corollary of Proposition 2.4 and Proposition 2.12 of [20]. To prove Lemma 2, we first need to show that there exists uniformly most powerful test for the problem:
$H_{0}$: the state is $s_{0}$. versus $H_{1}$: the state is $s_{1}$.
It is straightforward to see that this is a simple binary hypothesis testing problem. Hence, Neyman-Pearson lemma implies that the likelihood-ratio test is the UMP test. Therefore, for any significance level $\alpha\in[0,1]$, there is
$\beta^{\alpha}=\inf\nolimits_{\phi}\left\{\beta_{\phi}:\alpha_{\phi}\leq\alpha\right\},$
such that $\alpha=\mathbb{E}[\phi|H_{0},\gamma]$ and $\beta^{\alpha}=1-\mathbb{E}[\phi|H_{1},\gamma]$. Therefore, the mechanism $\mathcal{M}(\gamma)$ is $f$-differentially private (Definition 2.3 of [20]). Then, by Proposition 2.4 of [20], $\mathcal{M}(\gamma)$ is also $f^{*}$-differentially private, where $f^{*}(y)=\sup\nolimits_{0\leq\alpha^{\prime}\leq 1}y\alpha^{\prime}-f(\alpha^{\prime})$. Then, we obtain $\delta(\epsilon)=1+f^{*}(-e^{\epsilon})$ from Proposition 2.12 of [20].
Lemma 2 shows that any randomized mechanism with well-defined probability distributions satisfies $(\epsilon,\delta)$-DCP for some $\epsilon\geq 0$ and $\delta\in[0,1]$. It is a corollary of Proposition 2.4 and Proposition 2.12 of [20].
Lemma 3
Fix any $\theta$. Suppose that each $\mathcal{M}_{i}(\gamma_{i})$ is $(\epsilon_{i},\delta_{i})$-DCP with $\epsilon_{i}\geq 0$ and $\delta_{i}\in[0,1)$ for all $i\in[k]$, and $\mathcal{M}(\vec{\gamma})$ is the $k$-fold composition with $f_{[k]}$ as the optimal symmetric trade-off function. Then, the mechanism $\mathcal{M}(\vec{\gamma})$ is $(\epsilon^{\prime}_{g},\delta(\epsilon^{\prime}_{g}))$-DCP for some $\epsilon^{\prime}_{g}\geq 0$.
Proof:
Consider the following hypothesis testing: $H_{0}$: the state is $s_{0}$. versus $H_{1}$: the state is $s_{1}$. Under the composition, given any $\vec{y}$, the likelihood ratio is given by
$\textup{LR}(\vec{y})=\frac{\ell(\vec{y}|H_{0})}{\ell(\vec{y}|H_{1})},$
where
$\ell(\vec{y}|H_{z})=\int\limits_{x}\prod_{i=1}^{k}\gamma_{i}(y_{i}|x)\frac{P_{\theta}(s_{z},x)}{P_{\theta}(s_{z})}dx,$
for $z\in\{0,1\}$, which is well-defined. Then, by Neyman-Pearson lemma, the likelihood-ratio test is the UMP test and the proof follows Lemma 2.
A-G Proof of Parts (ii) and (iii) of Theorem 1 and Theorem 3
For any $\theta\in\Theta$, we define $\mathcal{Q}\equiv\{(s,s^{\prime})\mid\mathtt{D}(s,s^{\prime}\mid\theta)\leq\mathtt{d}\}$ for simplicity. The proof of this theorem involves notions and definitions from hypothesis testing, as outlined in Appendix A-F1. We start by proving part (i) of Theorem 3.
A-G1 Proof of Part (i) of Theorem 3
When $\textbf{L}^{\mathcal{G}}_{c_{s_{0}},c_{s_{1}}}(\vec{Y})=0$ for all $(s_{0},s_{1})\in\mathcal{Q}$, the PLRV under composition becomes $\textbf{L}^{\textup{id}}_{\psi_{s_{0}},\psi_{s_{1}}}(\vec{Y})$, the sum of independent PLRVs. In addition, the basic composition of DP implies that the composition $\mathcal{M}(\vec{\gamma})$ is $(\hat{\epsilon}_{g},\delta^{\dagger}_{g})$-DCP, where $\hat{\epsilon}_{g}=\sum^{k}_{i=1}\epsilon_{i}$ and $\delta^{\dagger}_{g}=\sum^{k}_{i=1}\delta_{i}$.
Let
$\hat{\delta}_{g}\equiv\mathbb{E}_{\textbf{L}\sim\textup{PLD}^{\mathtt{in}}}\left[\left(1-e^{\hat{\epsilon}_{g}-\textbf{L}}\right)^{+}\right],$
where $\textup{PLD}^{\mathtt{in}}$ is the PLD of $\textbf{L}^{\textup{id}}_{\psi_{s_{0}},\psi_{s_{1}}}(\vec{Y})$. In addition, let $\epsilon^{\dagger}_{g}$ such that
$\mathbb{E}_{\textbf{L}\sim\textup{PLD}^{\mathtt{in}}}\left[\left(1-e^{\epsilon^{\dagger}_{g}-\textbf{L}}\right)^{+}\right]=\delta^{\dagger}_{g}\equiv\sum^{k}_{i=1}\delta_{i}.$
By Proposition 7 of [36], (when $\textbf{L}^{\mathcal{G}}_{c_{s_{0}},c_{s_{1}}}(\vec{Y})=0$), the following holds:
  • If we fix the ratio at $e^{\hat{\epsilon}_{g}}$, the additive ‘excess’ that the event’s probability under $s$ can have—beyond that scaled probability under $s^{\prime}$—is at most $\hat{\delta}_{g}$, where $\hat{\delta}_{g}\leq\delta^{\dagger}_{g}$.
  • If we fix the additive $\delta^{\dagger}_{g}$, there is no attacker that can violate privacy in terms of epsilon-delta indistinguishability by more than $e^{\epsilon^{\dagger}_{g}}\leq e^{\hat{\epsilon}_{g}}$.
For any $D>0$, define
$F(D;\epsilon)\equiv\mathbb{E}_{\textbf{L}\sim\textup{PLD}^{\mathtt{in}}}\left[\left(1-\frac{e^{\epsilon-\textbf{L}}}{D}\right)^{+}\right].$
For any fixed $\epsilon\geq 0$, it is easy to verify that $F(D;\epsilon)$ is a continuous and strictly increasing function of $D$ for all $D>0$. In addition, the Intermediate Value Theorem ensures that for any $W\in(0,1]$, there is a $D$ such that $F(D;\epsilon)=W$ for any fixed $\epsilon\geq 0$. Furthermore, for any $\epsilon\geq 0$, there is a $D$ such that $F(D;\epsilon)=W$ for any fixed $W\in(0,1]$.
Consider:
  • Let $\hat{D}>1$ such that when $\hat{\epsilon}_{g}$ is fixed, $F(1;\hat{\epsilon}_{g})$ is increased to $F(\hat{D};\hat{\epsilon}_{g})=\delta^{\dagger}_{g}$.
  • Let $D^{\dagger}>0$ such that when $\delta^{\dagger}_{g}$ is fixed, $\epsilon^{\dagger}$ is changed to $\hat{\epsilon}_{g}$ such that $\delta^{\dagger}_{g}=F(D^{\dagger};\hat{\epsilon}_{g})$.
Now, suppose that $\textbf{L}^{\mathcal{G}}_{c_{s_{0}},c_{s_{1}}}(\vec{Y})\neq 0$ in general. Then, we have, for any $\epsilon\geq 0$,
$\mathbb{E}_{\textbf{L}\sim\textup{PLD}}\left[\left(1-\frac{e^{\epsilon-\textbf{L}^{\mathtt{in}}}}{e^{\textbf{L}^{\mathcal{G}}}}\right)^{+}\right]=F(e^{\textbf{L}^{\mathcal{G}}};\epsilon).$
The basic composition theory holds for DCP when $\textbf{L}^{\mathcal{G}}_{c_{s_{0}},c_{s_{1}}}(\vec{Y})\neq 0$ if and only if
  • For given $\hat{\epsilon}_{g}=\sum^{k}_{i=1}\epsilon_{i}$, the basic composition theory holds for DCP if and only if $e^{\textbf{L}^{\mathcal{G}}}\leq\hat{D}$ or $\textbf{L}^{\mathcal{G}}\leq\log\hat{D}$.
  • For given $\delta^{\dagger}_{g}=\sum^{k}_{i=1}\delta_{i}$, the basic composition theory holds for DCP if and only if $e^{\textbf{L}^{\mathcal{G}}}\leq D^{\dagger}$ or $\textbf{L}^{\mathcal{G}}\leq\log D^{\dagger}$.
However, $\textbf{L}^{\mathcal{G}}_{c_{s_{0}},c_{s_{1}}}(\vec{Y})\neq 0$, in general, cannot guarantee $\textbf{L}^{\mathcal{G}}\leq\log\hat{D}$ or $\textbf{L}^{\mathcal{G}}\leq\log D^{\dagger}$.
A-G2 Proof of Part (ii) of Theorem 3
The following theorem aligns with part (ii) of Theorem 3.
Theorem 5
Fix any $\theta$. Suppose that each $\mathcal{M}_{i}(\gamma_{i})$ is $(\epsilon_{i},\delta_{i})$-CP with $\epsilon_{i}\geq 0$ and $\delta_{i}\in[0,1)$ for all $i\in[k]$, and let $\mathcal{M}(\vec{\gamma})$ be the composition mechanism with $f_{[k]}$ as the optimal symmetric trade-off function. Let
$\epsilon^{*}(\delta_{g})=\textup{OPT}\left(\mathcal{N}(\psi_{[k]}),\delta_{g}\right),$
for any $\delta_{g}\in[0,1)$, where $\vec{\psi}$ is induced by $\vec{\gamma}$ and OPT is given by (5). Additionally, let
$\delta(\epsilon_{g})=1+f^{*}_{[k]}(-e^{\epsilon_{g}}),$
for any $\epsilon_{g}\geq 0$. Then, the following holds:
  • (i)
    For any $\epsilon_{g}\geq 0$ such that $\delta(\epsilon_{g})\in[0,1)$, we have $\epsilon_{g}\geq\epsilon^{*}\left(\delta(\epsilon_{g})\right)$.
  • (ii)
    For any $\delta_{g}\in[0,1)$, $\delta_{g}\geq\delta(\epsilon^{*}(\delta_{g}))$.
Equality holds in (i) and (ii) if and only if
$\vec{\psi}(\vec{y}\mid s)=\prod_{i=1}^{k}\psi_{i}(y_{i}\mid s),$
for all $s\in\mathcal{S}$, $\vec{y}\in\prod_{i=1}^{k}\mathcal{Y}_{i}$, and $i\in[k]$.
Theorem 5 demonstrates that the optimal bound on privacy loss, as captured by the optimal composition of standard differential privacy [6], underestimates the privacy degradation in the composition of CP mechanisms, even when the individual mechanisms are independent of each other.
Both Theorems 5 and 3 illustrate that the tightest privacy bound under the optimal composition of DP underestimates the actual cumulative privacy loss of CP. We prove these two theorems together in this section.
For simplicity, we restrict our attention to the composition of two individual mechanisms, without loss of generality. Consider two mechanisms $\mathcal{M}_{1}(\gamma_{1})$ and $\mathcal{M}_{2}(\gamma_{2})$ that are $(\epsilon_{1},\delta_{1})$-CP and $(\epsilon_{2},\delta_{2})$-CP, respectively. Let $\mathcal{N}_{i}(\psi_{i})$ be induced by $\mathcal{M}_{i}(\gamma_{i})$ for $i\in\{1,2\}$.
Recall that the density function $\psi_{i}$ is given by
[…]
which is induced by $\gamma_{i}$ and $\theta$.
Define
$\mathtt{D}_{1,2}(y_{1},y_{2}\mid s)=\psi_{1}(y_{1}\mid s)\psi_{2}(y_{2}\mid s).$
Additionally, the density function associated with $\mathcal{M}(\vec{\gamma})$ is given by
$\mathtt{D}_{[2]}(y_{1},y_{2}\mid s)=\int_{x}\prod_{i=1}^{2}\gamma_{i}(y_{i}\mid x)\frac{P_{\theta}(s,x)}{P_{\theta_{S}}(s)}dx.$
We adopt the general notation of the trade-off function as given in Definition 2.1 of [20].
Definition 8 (Trade-off Function)
Let $\alpha_{\phi}$ and $\beta_{\phi}$ be the Type-I and Type-II errors associated with any rejection rule $\phi$, respectively. For any two probability distributions $P$ and $Q$ on the same space, the trade-off function $T(P,Q):[0,1]\to[0,1]$ is defined as
$T(P,Q)(\alpha)=\inf\{\beta_{\phi}:\alpha_{\phi}\leq\alpha\},$
where the infimum is taken over all (measurable) rejection rules.
The following Proposition 4 shows that for any $\vec{\gamma}$, $\mathtt{D}_{[2]}$ is more informative than $\mathtt{D}_{1,2}$ about the state in the sense of Blackwell [38, 39].
Proposition 4
For any $(s_{1},s_{0})\in\mathcal{Q}$ and $\alpha\in[0,1]$, we have
$T(\mathtt{D}_{[2]}(\cdot\mid s_{0}),\mathtt{D}_{[2]}(\cdot\mid s_{1}))\leq T(\mathtt{D}_{1,2}(\cdot\mid s_{0}),\mathtt{D}_{1,2}(\cdot\mid s_{1})),$
where equality holds if and only if $\mathtt{D}_{[2]}=\mathtt{D}_{1,2}$.
The proof of Proposition 4 relies on Lemmas 4 and 5. Specifically:
  • Lemma 4 shows that Bayesian inference based on $\mathtt{D}_{[2]}$ results in a smaller expected cross-entropy loss compared to that based on $\mathtt{D}_{1,2}$.
  • Lemma 5 establishes an equivalence between the ordering of expected cross-entropy losses and Blackwell’s ordering of informativeness.
Thus, by applying Lemmas 4 and 5, we can prove Proposition 4.
Detailed Proofs
By Bayes’ rule, the posterior distributions are constructed as follows:
$\textup{J}_{1,2}(s\mid y_{1},y_{2})=\frac{\mathtt{D}_{1,2}(y_{1},y_{2}\mid s)p_{\theta_{S}}(s)}{\textup{P}_{1,2}(y_{1},y_{2})},$
$\textup{J}_{[2]}(s\mid y_{1},y_{2})=\frac{\mathtt{D}_{[2]}(y_{1},y_{2}\mid s)p_{\theta_{S}}(s)}{\textup{P}_{[2]}(y_{1},y_{2})},$
[…]
For any randomized inference strategy $\sigma:\mathcal{Y}_{1}\times\mathcal{Y}_{2}\to\Delta(\mathcal{S})$, the expected cross-entropy loss (CEL) induced by $\{\gamma_{1},\gamma_{2},\theta\}$ is given by
$\textup{CEL}(\sigma)=$ $-\int_{s,x}p_{\theta}(s,x)\int_{y_{1},y_{2}}\prod_{i=1}^{2}\gamma_{i}(y_{i}\mid x)$
×$\times\log\sigma(s\mid y_{1},y_{2})\,dx\,ds\,dy_{[2]}.$
Lemma 4
The following inequality holds:
$\textup{CEL}\left(\textup{J}_{[2]}\right)\leq\textup{CEL}\left(\textup{J}_{1,2}\right),$
with equality if and only if $\textup{J}_{[2]}=\textup{J}_{1,2}$.
Proof:
We consider the difference between the expected CELs:
$\textup{CEL}\left(\textup{J}_{1,2}\right)-\textup{CEL}\left(\textup{J}_{[2]}\right)$ $=-\int_{s,x}p_{\theta}(s,x)\int_{y_{1},y_{2}}\prod_{i=1}^{2}\gamma_{i}(y_{i}\mid x)$
×$\quad\times\log\frac{\textup{J}_{[2]}(s\mid y_{1},y_{2})}{\textup{J}_{1,2}(s\mid y_{1},y_{2})}\,dx\,ds\,dy_{[2]}.$
By the definitions of $\mathtt{D}_{[2]}$ and $\textup{J}_{[2]}$, we have:
$\textup{J}_{[2]}(s\mid y_{1},y_{2})\textup{P}_{[2]}(y_{1},y_{2})=\int_{x}\prod_{i=1}^{2}\gamma_{i}(y_{i}\mid x)p_{\theta}(s,x)\,dx.$
Using this, the difference $\textup{CEL}\left(\textup{J}_{1,2}\right)-\textup{CEL}\left(\textup{J}_{[2]}\right)$ can be rewritten as the Kullback–Leibler (KL) divergence between $\textup{J}_{[2]}$ and $\textup{J}_{1,2}$:
$\textup{CEL}\left(\textup{J}_{1,2}\right)-\textup{CEL}\left(\textup{J}_{[2]}\right)$ $=\textup{KL}\left(\textup{J}_{[2]}\|\textup{J}_{1,2}\right),$
where
$\textup{KL}\left(\textup{J}_{[2]}\|\textup{J}_{1,2}\right)$ $=\int_{s}\int_{y_{1},y_{2}}\textup{J}_{[2]}(s\mid y_{1},y_{2})$
×$\quad\times\log\frac{\textup{J}_{[2]}(s\mid y_{1},y_{2})}{\textup{J}_{1,2}(s\mid y_{1},y_{2})}\textup{P}_{[2]}(y_{1},y_{2})\,dy_{1}\,dy_{2}\,ds.$
Since $\textup{KL}\left(\textup{J}_{[2]}\|\textup{J}_{1,2}\right)\geq 0$, with equality if and only if $\textup{J}_{[2]}=\textup{J}_{1,2}$, it follows that:
$\textup{CEL}\left(\textup{J}_{[2]}\right)\leq\textup{CEL}\left(\textup{J}_{1,2}\right).$
Lemma 5
Let $\mathcal{A}(\sigma)$ with $\sigma:\mathcal{Y}_{1}\times\mathcal{Y}_{2}\to\Delta(\mathcal{S})$ be any attack model. Let $\textup{J}_{[2]}$ and $\textup{J}_{1,2}$ be the posterior distributions induced by $\mathtt{D}_{[2]}$ and $\mathtt{D}_{1,2}$, respectively. Then, the following two statements are equivalent:
  • (i)
    $\textup{CEL}\left(\textup{J}_{[2]}\right)\leq\textup{CEL}\left(\textup{J}_{1,2}\right)$.
  • (ii)
    For any $(s_{1},s_{0})\in\mathcal{Q}$ and $\alpha\in[0,1]$,
    $T(\mathtt{D}_{[2]}(\cdot\mid s_{0}),\mathtt{D}_{[2]}(\cdot\mid s_{1}))\leq T(\mathtt{D}_{1,2]}(\cdot\mid s_{0}),\mathtt{D}_{1,2}(\cdot\mid s_{1})).$
Equality holds if and only if $\mathtt{D}_{[2]}=\mathtt{D}_{1,2}$.
Proof:
By Theorem 1 of [39], part (ii) of Lemma 5 is equivalent to the following:
  • There exists a randomized mechanism $\mathtt{Proc}:\mathcal{Y}_{1}\times\mathcal{Y}_{2}\to\mathcal{Z}$ such that
    $\mathtt{D}_{1,2}(s_{0})=\mathtt{Proc}(\mathtt{D}_{[2]}(s_{0})),\;\mathtt{D}_{1,2}(s_{1})=\mathtt{Proc}(\mathtt{D}_{[2]}(s_{1})),$
    for all $(s_{0},s_{1})\in\mathcal{Q}$.
Then, by Theorem 10 in [38] (see also Theorem 2.10 in [20] and Theorem 2.5 in [5]), we establish the equivalence of parts (i) and (ii) of Lemma 5. ∎
Recall that a function $f:[0,1]\to[0,1]$ is a symmetric trade-off function if it is convex, continuous, non-increasing, satisfies $f(\alpha)\leq 1-\alpha$ and
$f(\alpha)=\inf\{t\in[0,1]:f(t)\leq\alpha\},$
for $\alpha\in[0,1]$ [20].
The convex conjugate of $f$ is defined as
$f^{*}(y)=\sup\nolimits_{0\leq\alpha^{\prime}\leq 1}\,(y\alpha^{\prime}-f(\alpha^{\prime})).$
Let $\epsilon^{*}(\delta_{g})=\textup{OPT}\left(\mathcal{N}(\psi_{[2]}),\delta_{g}\right)$ for any $\delta_{g}\in[0,1)$, where $\psi_{[2]}$ is induced by $\gamma_{[2]}$. Additionally, define
$\delta_{[2]}(\epsilon_{g})=1+f^{*}_{[2]}(-e^{\epsilon_{g}}),$
$\delta_{1,2}(\epsilon_{g})=1+f^{*}_{1,2}(-e^{\epsilon_{g}}),$
for any $\epsilon_{g}\geq 0$.
When $\mathtt{D}_{[2]}\neq\mathtt{D}_{1,2}$, Proposition 4 implies
$T(\mathtt{D}_{[2]}(\cdot\mid s_{0}),\mathtt{D}_{[2]}(\cdot\mid s_{1}))<T(\mathtt{D}_{1,2}(\cdot\mid s_{0}),\mathtt{D}_{1,2}(\cdot\mid s_{1})).$
Consequently, $f_{[2]}(\alpha)<f_{1,2}(\alpha)$ for all $\alpha\in[0,1]$, which implies that for any fixed $\epsilon_{g}$ and $\epsilon$, we have $\delta_{[2]}(\epsilon)>\delta_{1,2}(\epsilon)$.
Suppose $\delta_{[2]}(\epsilon_{[2]})=\delta_{1,2}(\epsilon_{1,2})$. Then, $f^{*}_{[2]}(-e^{\epsilon^{[2]}_{g}})=f^{*}_{1,2}(-e^{\epsilon^{1,2}_{g}})$. Since both $f_{[2]}$ and $f_{1,2}$ are convex, continuous, and non-increasing functions, their convex conjugates are concave, continuous, and non-decreasing functions.
Furthermore, since $\delta_{[2]}(\epsilon)>\delta_{1,2}(\epsilon)$ for all $\epsilon\geq 0$, to achieve $\delta_{[2]}(\epsilon_{[2]})=\delta_{1,2}(\epsilon_{1,2})$, we must have $\epsilon_{[2]}>\epsilon_{1,2}$.
Thus, we complete the proof of Theorem 5.
Theorem 5 confirms the following parts of Theorem 3:
  • If $\textbf{L}_{c_{s_{0}},c_{s_{1}}}(\vec{Y})\neq 0$, then
    $\underline{\textbf{opt}}_{\delta_{g}}<\textbf{opt}_{\delta_{g}},\;\underline{\textbf{dt}}_{\epsilon_{g}}<\textbf{dt}_{\epsilon_{g}}.$
The second set of inequalities:
$\textbf{opt}_{\delta_{g}}\leq\overline{\textbf{opt}}_{\delta_{g}},\;\textbf{dt}_{\epsilon_{g}}\leq\overline{\textbf{dt}}_{\epsilon_{g}},$
arise from the fact that, in the PLRV given by (11), the copula-dependent PLRV $\textbf{L}^{\mathcal{G}}_{c_{s_{0}},c_{s_{1}}}(\vec{Y})$ is correlated with $\{\textbf{L}_{\psi_{i}^{s_{0}},\psi_{i}^{s_{1}}}(Y_{i})\}_{i=1}^{k}$, where each element of $\{\textbf{L}_{\psi_{i}^{s_{0}},\psi_{i}^{s_{1}}}(Y_{i})\}_{i=1}^{k}$ is independent of the others.
Consequently, even when each $\psi_{i}$ achieves the bounds of its probabilistic indistinguishability, the copula may not attain its own bound. The maximum privacy loss ($\overline{\textbf{opt}}_{\delta_{g}}$ or $\overline{\textbf{dt}}^{\epsilon_{g}}$) occurs when all the individual mechanisms and the copula term achieve their respective bounds of probabilistic indistinguishability, maximizing the cumulative privacy loss.
$\square$
A-H Proof of Theorem 2
In part (i), the condition $\Pi[\epsilon_{g},\delta_{g}]\neq\emptyset$ is necessary because the target privacy budget $(\epsilon_{g},\delta_{g})$ may already be violated by the existing $k$ mechanisms $\{\mathcal{M}_{i}(\gamma_{i})\}_{i=1}^{k}$.
Part (ii) of Theorem 1 (and part (i) of Theorem 3) indicates that while basic composition theory is generally satisfied by DCP, it prevents directly using a conservative summation of privacy parameters to estimate the overall privacy budget. Consequently, Task-1 requires a rough and conservative estimation of the total privacy loss to ensure that $\Pi[\epsilon_{g},\delta_{g}]\neq\emptyset$.
In Task-2, $\epsilon_{g}$ becomes an additional decision variable. Ideally, if the optimization problem is solved perfectly, $\epsilon_{g}$ would correspond to the true, tightest privacy loss value for the composition $\mathcal{M}(\vec{\gamma})$.
The main proof task of Theorem 2 is to demonstrate that the constraints specified by $\Pi[\epsilon_{g},\delta_{g}]$, for $\epsilon_{g}\geq 0$ and $\delta_{g}\in[0,1]$, are sufficient to guarantee that the privacy strategy $\alpha$—which induces the best response $\pi$ satisfying $\Pi[\epsilon_{g},\delta_{g}]$—ensures that the composition satisfies $(\epsilon_{g},\delta_{g})$-DCP.
Given any choice of strictly proper scoring rules $\mathcal{L}(\pi,s,\vec{y}_{\alpha})$, the posterior distribution $\pi^{*}=\mu$ is the unique minimizer of $\mathbb{E}_{\alpha,\vec{\gamma}}^{\mu}[\mathcal{L}(\pi,s,\vec{y}_{\alpha})]$ [31]. Here, $\mu$ denotes the posterior distribution induced by the leader’s privacy strategy $\alpha$ and the composition $\mathcal{M}(\vec{\gamma})$ (or equivalently, $\mathcal{M}(\vec{\gamma}_{\alpha})$).
To establish that the composition $\mathcal{M}(\vec{\gamma}_{\alpha})$ satisfies $(\epsilon_{g},\delta_{g})$-DCP, we demonstrate that the induced posterior distribution fulfills the conditions in $\Pi[\epsilon_{g},\delta_{g}]$. Consequently, the composition $\mathcal{M}(\vec{\gamma}_{\alpha})$ is proven to be $(\epsilon_{g},\delta_{g})$-DCP.
With abuse of notation, we let $\vec{\gamma}^{c}_{\alpha}$ denote the joint density function of $\mathcal{M}(\vec{\gamma}_{\alpha})$ with the underlying copula density $c$ capturing the dependencies among the mechanisms $\mathcal{M}_{1}(\gamma_{1}),\dots,\mathcal{M}_{k}(\gamma_{k})$. In addition, let $\vec{\psi}^{c}_{\alpha}(\cdot|s)$ be the corresponding density of the effective mechanism. Given the density $\vec{\gamma}^{c}_{\alpha}$ and any output profile $\vec{y}_{\alpha}\in\vec{\mathcal{Y}}\times\mathcal{Y}_{\alpha}$, the posterior belief is constructed by
$\text{J}\left(s\middle|\vec{y}_{\alpha}\right)=\frac{P(s,\vec{y}_{\alpha})}{P(s,\vec{y}_{\alpha})},$
[…]
By Lemma 6 shown below, we first obtain that
$\text{J}\left(s\middle|\vec{y}_{\alpha}\right)\geq e^{-\epsilon}P_{\theta}(s),$
and
$\mathbb{E}_{\text{J}}\left[\frac{\text{J}\left(S\middle|\vec{y}_{\alpha}\right)}{P_{\theta}(S)}\right]\leq\delta_{g}e^{\epsilon_{g}},$
implies
$\text{Pr}_{S\sim\text{J}\left(\cdot\middle|\vec{y}_{\alpha}\right)}\left[e^{-\epsilon_{g}}\leq\frac{\text{J}\left(S\middle|\vec{y}_{\alpha}\right)}{P_{\theta}(S)}\leq e^{\epsilon_{g}}\right]\geq 1-\delta_{g},$
for all $\vec{y}_{\alpha}\in\vec{\mathcal{Y}}\times\mathcal{Y}_{\alpha}$. Then, if $\vec{Y}_{\alpha}\sim\vec{\psi}^{c}_{\alpha}(\cdot|s)$, it holds that
$\text{Pr}_{\vec{Y}_{\alpha}\sim\vec{\psi}^{c}_{\alpha}(\cdot|s)}\left[e^{-\epsilon_{g}}\leq\frac{\text{J}\left(S\middle|\vec{Y}_{\alpha}\right)}{P_{\theta}(S)}\leq e^{\epsilon_{g}}\right]\geq 1-\delta_{g}.$ (13)
From Lemma 7, we have that
$e^{-\epsilon_{g}}\leq\frac{\text{J}\left(s\middle|\vec{y}_{\alpha}\right)}{P_{\theta}(s)}\leq e^{\epsilon_{g}},\forall s\in\mathcal{S},$
implies
$\textbf{Pr}^{s}_{\vec{\gamma}^{c}_{\alpha}}\left[\vec{y}_{\alpha}\in\vec{\mathcal{W}}_{\alpha}\right]\leq e^{\varepsilon(\epsilon_{g},\theta)}\textbf{Pr}^{s^{\prime}}_{\vec{\gamma}^{c}_{\alpha}}\left[\vec{y}_{\alpha}\in\vec{\mathcal{W}}_{\alpha}\right],$
for all $(s,s^{\prime})\in\mathcal{Q}$, $\vec{\mathcal{W}}_{\alpha}\subset\vec{\mathcal{Y}}_{\alpha}$, where $\varepsilon(\epsilon_{g},\theta)\equiv\log\left(1+\frac{e^{\epsilon_{g}}-1}{P^{*}_{\theta}}\right)$ with $P^{*}_{\theta}=\min_{s}P_{\theta}(s)$. Thus, by (13), we have
$\text{Pr}_{\vec{Y}_{\alpha}\sim\vec{\psi}^{c}_{\alpha}(\cdot|s)}\left[\frac{\textbf{Pr}^{s}_{\vec{\gamma}^{c}_{\alpha}}\left[\vec{y}_{\alpha}\in\vec{\mathcal{W}}_{\alpha}\right]}{\textbf{Pr}^{s^{\prime}}_{\vec{\gamma}^{c}_{\alpha}}\left[\vec{y}_{\alpha}\in\vec{\mathcal{W}}_{\alpha}\right]}\leq e^{\epsilon_{g}}\right]\geq 1-\delta_{g},$ (14)
for all $(s,s^{\prime})\in\mathcal{Q}$, $\vec{\mathcal{W}}_{\alpha}\subset\vec{\mathcal{Y}}_{\alpha}$.
Finally, Lemma 8 shows that (14) implies
$\textbf{Pr}^{s}_{\vec{\gamma}^{c}_{\alpha}}\left[\vec{y}_{\alpha}\in\vec{\mathcal{W}}_{\alpha}\right]\leq e^{\epsilon_{g}}\textbf{Pr}^{s^{\prime}}_{\vec{\gamma}^{c}_{\alpha}}\left[\vec{y}_{\alpha}\in\vec{\mathcal{W}}_{\alpha}\right]+\delta_{g}.$
Therefore, we can conclude that the sufficiency of the conditions given by $\Pi[\epsilon_{g},\delta_{g}]$ for any $\epsilon_{g}\geq 0$ and $\delta_{g}\in(0,1]$. In addition, when $\delta_{g}=0$, Lemma 7 establishes the sufficiency.
A-H1 Related Lemmas
Lemma 6
Let $F:\mathcal{Y}\mapsto\Delta(\mathcal{S})$ be any conditional probability mass or density function, and let $P_{S}\in\Delta(\mathcal{S})$ be any prior distribution of the secret $S$. For any $\epsilon\geq 0$ and $\delta\in[0,1)$, if
$F(s|y)\geq e^{-\epsilon}p_{\theta_{S}}(s)\textup{ and }\mathbb{E}\left[\frac{F\left(S\middle|y\right)}{P_{S}(S)}\right]\leq\delta e^{\epsilon},$
where the expectation is taken over the randomness of $S\sim F(\cdot|y)$, then
$\textup{Pr}\left[e^{-\epsilon}\leq\frac{F(S|y)}{P_{S}(S)}\leq e^{\epsilon}\right]\geq 1-\delta,$
for any $s\in\{s_{0},s_{1}\}\in\mathcal{Q}$, $y\in\mathcal{Y}$.
Proof:
First, $F(s|y)\geq e^{-\epsilon}p_{\theta_{S}}(s)$ implies
$\textup{Pr}\left[e^{-\epsilon}\leq\frac{F(s|y)}{P_{S}(s)}\leq e^{\epsilon}\right]=1-\textup{Pr}\left[\frac{F(s|\vec{y}_{[k]})}{P_{S}(s)}\geq e^{\epsilon}\right].$
By Markov’s inequality, we have
$\textup{Pr}\left[\frac{F(s|\vec{y}_{[k]})}{P_{S}(s)}\geq e^{\epsilon}\right]\leq\frac{\mathbb{E}\left[\frac{F(s|\vec{y}_{[k]})}{P_{S}(s)}\right]}{e^{\epsilon}}.$
Since $\mathbb{E}\left[\frac{F\left(S\middle|y\right)}{P_{S}(S)}\right]\leq\delta e^{\epsilon}$, we have
$\textup{Pr}\left[\frac{F(s|\vec{y}_{[k]})}{P_{S}(s)}\geq e^{\epsilon}\right]\leq\delta.$
Therefore, $\textup{Pr}\left[e^{-\epsilon}\leq\frac{F(s|\vec{y}_{[k]})}{P_{S}(s)}\leq e^{\epsilon}\right]\geq 1-\delta$. ∎
Lemma 7
Given $\theta\in\Theta$, a mechanism $\mathcal{M}(\gamma)$ is $(\varepsilon(\epsilon_{g},\theta),0)$-DCP with $\varepsilon(\epsilon_{g},\theta)=\log\left(1+\frac{e^{\epsilon_{g}}-1}{P^{*}_{\theta}}\right)$ where $P^{*}_{\theta}=\min_{s}P_{\theta}(s)$ if the posterior J satisfies
$e^{-\epsilon}\leq\frac{\text{J}\left(s\middle|y\right)}{P_{\theta}(s)}\leq e^{\epsilon},$ (15)
for all $s\in\mathcal{S}$.
Lemma 7 extends Claim 3 of [1] to the general differential confounding privacy.
Proof:
Let $\mathcal{N}(\psi):\mathcal{S}\mapsto\mathcal{Y}$ be the effective mechanism of $\mathcal{M}(\gamma)$, with well-defined density $\psi$, prior $P_{\theta}$, and induced posterior $\mathtt{J}$. Let $E$ be any measurable event.
Suppose that the mechanism $\mathcal{M}(\gamma)$ satisfies (15). Then,
$e^{-\epsilon}\leq\frac{\text{J}\left(E\middle|y\right)}{P_{\theta}(E)}\leq e^{\epsilon}.$
Given $\theta\in\Theta$, construct a prior $P_{\theta}$ (with abuse of notation) that places mass only on $s_{0}$ and $s_{1}$, such that $P_{\theta}(s_{0})+P_{\theta}(s_{1})=1$. Set the event $E=\{s_{1}\}$. Then,
$\mathtt{J}(E|y)=\frac{\psi(y|s_{1})P_{\theta}(s_{1})}{\psi(y|s_{0})P_{\theta}(s_{0})+\psi(y|s_{1})P_{\theta}(s_{1})},$
which gives
$\frac{P_{\theta}(E)}{\text{J}(E|y)}$ $=P_{\theta}(s_{1})+P_{\theta}(s_{0})\frac{\psi(y|s_{0})}{\psi(y|s_{1})}$
$=1-P_{\theta}(s_{0})+P_{\theta}(s_{0})\frac{\psi(y|s_{0})}{\psi(y|s_{1})}.$
Since (15) is satisfied, we have
$e^{-\epsilon}\leq 1-P_{\theta}(s_{0})+P_{\theta}(s_{0})\frac{\psi(y|s_{0})}{\psi(y|s_{1})}\leq e^{\epsilon},$
which yields
$\max\{0,\frac{e^{-\epsilon}-1}{P_{\theta}(s_{0})}+1\}\leq\frac{\psi(y|s_{0})}{\psi(y|s_{1})}\leq\frac{e^{\epsilon}-1}{P_{\theta}(s_{0})}+1,$
for all $(s_{0},s_{1})\in\mathcal{Q}$. Therefore, the maximum upper bound of the ratio $\frac{\psi(y|s_{0})}{\psi(y|s_{1})}$, for all $(s_{0},s_{1})\in\mathcal{Q}$, that we can have is $\varepsilon(\epsilon_{g},\theta)=\log\left(1+\frac{e^{\epsilon_{g}}-1}{\min_{s}P_{\theta}(s)}\right)$, where $\min_{s}P_{\theta}(s)>0$ for all $s\in\{s_{0},s_{1}\}\in\mathcal{Q}$.
The condition (16) in Lemma 8 is equivalent to the strong Bayesian DP introduced by [37]. By treating the DCP as a single-point DP, Lemma 8 can be proved in the similar way as Proposition 1 of [37].
Lemma 8
Given $\theta\in\Theta$, if
$\text{Pr}_{\vec{Y}_{\alpha}\sim\vec{\psi}^{c}_{\alpha}(\cdot|s)}\left[\frac{\textbf{Pr}^{s}_{\vec{\gamma}^{c}_{\alpha}}\left[\vec{y}_{\alpha}\in\vec{\mathcal{W}}_{\alpha}\right]}{\textbf{Pr}^{s^{\prime}}_{\vec{\gamma}^{c}_{\alpha}}\left[\vec{y}_{\alpha}\in\vec{\mathcal{W}}_{\alpha}\right]}\leq e^{\epsilon_{g}}\right]\geq 1-\delta_{g},$ (16)
for all $\vec{\mathcal{W}}_{\alpha}\subseteq\vec{\mathcal{Y}}\times\mathcal{Y}_{\alpha}$, $(s,s^{\prime})\in\mathcal{Q}$, $\epsilon_{g}\geq 0$, $\delta_{g}\in[0,1)$, then we have
$\textbf{Pr}^{s}_{\vec{\gamma}^{c}_{\alpha}}\left[\vec{y}_{\alpha}\in\vec{\mathcal{W}}_{\alpha}\right]\leq e^{\epsilon_{g}}\textbf{Pr}^{s^{\prime}}_{\vec{\gamma}^{c}_{\alpha}}\left[\vec{y}_{\alpha}\in\vec{\mathcal{W}}_{\alpha}\right]+\delta_{g}.$
A-I Experimental Details
A-I1 Network Configurations and Hyperparameters
The Defender neural network is a generative model designed to process membership vectors and generate beacon modification decisions. The input layer is followed by a series of fully connected layers with activation functions applied after each layer. The first and second hidden layers use ReLU activation, while the third hidden layer employs LeakyReLU activation. Layer normalization is applied after the second and third hidden layers. The output layer utilizes a custom ScaledSigmoid activation function, producing a real value in a scaled range determined by this function. The network is trained using the AdamW optimizer with a learning rate of $1.5\times 10^{-5}$ and weight decay of $1.5\times 10^{-5}$. Specific configurations for Defender performing single-shot re-design and copula perturbation are provided in Tables IV and V, respectively.
The Attacker neural network is a generative model designed to process beacons and noise, producing membership vectors. The input layer is followed by multiple fully connected layers, each equipped with batch normalization and ReLU activation. The output layer applies a standard sigmoid activation function. All Attacker models were trained using the AdamW optimizer, with a learning rate of $1.0\times 10^{-4}$ and a weight decay of $1.0\times 10^{-4}$. Specific configurations for Attacker when Defender performing single-shot re-design and copula perturbation are provided in Tables VI and VII, respectively.
Defender Input Units Output Units
Input Layer 800 5000
Hidden Layer 1 5000 9000
Hidden Layer 2 9000 3000
Output Layer 3000 1000
TABLE IV: Defender Neural Network Architecture (Single Re-Design)
Defender Input Units Output Units
Input Layer 830 3000
Hidden Layer 1 3000 2000
Hidden Layer 2 2000 1200
Hidden Layer 3 1200 1200
Output Layer 1200 1
TABLE V: Defender Neural Network Architecture (Copula Perturbation)
Attacker Input Units Output Units
Input Layer 1500 14000
Hidden Layer 1 14000 9500
Hidden Layer 2 9500 4500
Output Layer 4500 800
TABLE VI: Attacker Neural Network Architecture (Single Re-Design)
Attacker Input Units Output Units
Input Layer 1500 31000
Hidden Layer 1 31000 25000
Hidden Layer 2 25000 15000
Output Layer 15000 800
TABLE VII: Attacker Neural Network Architecture (Copula Perturbation)
A-I2 AUC Values with Standard Deviations
Tables VIII and IX show the AUC values shown in the plots.
Target &\&& Ind. $\epsilon$ Mechanisms AUC ± std
$\epsilon_{g}=0.25$, $\epsilon_{i}=0.05$ Composition 0.7441 ± 0.0055
Single 0.7485 ± 0.0064
$\epsilon_{g}=0.5$, $\epsilon_{i}=0.1$ Composition 0.8093 ± 0.0042
Single 0.8165 ± 0.0044
$\epsilon_{g}=1.5$, $\epsilon_{i}=0.3$ Composition 0.8921 ± 0.0031
Single 0.9056 ± 0.0029
$\epsilon_{g}=3$, $\epsilon_{i}=0.6$ Composition 0.9138 ± 0.0031
Single 0.9175 ± 0.0031
$\epsilon_{g}=5$, $\epsilon_{i}=1$ Composition 0.9330 ± 0.0029
Single 0.9308 ± 0.0028
TABLE VIII: AUC and standard deviation for single-shot re-design when the target $\delta_{g}$ and each individual $\delta_{i}$ are the same, i.e., $\delta_{g}=\delta_{i}=0.02$, and $\rho=0.5$
Target &\&& Ind. $\epsilon$ Type AUC ± std
$\epsilon_{g}=0.4$, $\epsilon_{i}=0.05$ Composition 0.5624 ± 0.0078
Single 0.5745 ± 0.0060
$\epsilon_{g}=0.6$, $\epsilon_{i}=0.1$ Composition 0.6206 ± 0.0071
Single 0.6388 ± 0.0053
$\epsilon_{g}=1$, $\epsilon_{i}=0.18$ Composition 0.6943 ± 0.0063
Single 0.7086 ± 0.0052
$\epsilon_{g}=2$, $\epsilon_{i}=0.3$ Composition 0.7429 ± 0.0055
Single 0.7324 ± 0.0051
$\epsilon_{g}=4$, $\epsilon_{i}=0.6$ Composition 0.8235 ± 0.0035
Single 0.8139 ± 0.0042
$\epsilon_{g}=6$, $\epsilon_{i}=1$ Composition 0.8238 ± 0.0043
Single 0.8142 ± 0.0044
TABLE IX: AUC and standard deviation for copula perturbation when the target $\delta_{g}$ and each individual $\delta_{i}$ are the same, i.e., $\delta_{g}=\delta_{i}=0.02$, and $\rho=0.5$